### v1

# Design Optimizer for Soft Growing Robot Manipulators in Three-Dimensional Environments

Abstract
Soft growing robots are novel devices that mimic plant-like growth for navigation in cluttered or dangerous environments. Their ability to adapt to surroundings, combined with advancements in actuation and manufacturing technologies, allows them to perform specialized manipulation tasks. This work presents an approach for design optimization of soft growing robots; specifically, the three-dimensional extension of the optimizer designed for planar manipulators. This tool is intended to be used by engineers and robot enthusiasts before manufacturing their robot: it suggests the optimal size of the robot for solving a specific task. The design process models a multi-objective optimization problem to refine a soft manipulator’s kinematic chain. Thanks to the novel Rank Partitioning algorithm integrated into Evolutionary Computation (EC) algorithms, this method achieves high precision in reaching targets and is efficient in resource usage. Results show significantly high performance in solving three-dimensional tasks, whereas comparative experiments indicate that the optimizer features robust output when tested with different EC algorithms, particularly genetic algorithms.
keywords:
Soft Robotics , Evolutionary Computation , Multi-Objective Optimization , Inverse Kinematics , Design Optimization , Rank Partitioning
PACS:
87.85.St , 87.55.de , 87.55.kd
MSC:
49
journal: ArXiv
[…]
1 Introduction
Soft robotics is a relatively new field of intelligent mechanical systems that aims at overcoming some of the biggest limitations of their rigid counterparts [1]. Rigid robots (i) may struggle to interact with fragile objects or delicate environments, such as human tissues, glassware, or sensitive components in archaeological sites; (ii) have intricate designs and their materials often result in high costs; and (iii) due to their strength and stiffness, the safety of human operators can be a concern in a human-machine interaction scenario [2]. On the other hand, soft robots are inspired by living organisms, with designs exploiting the natural compliance of their materials to facilitate smooth and intuitive physical interactions with the environment [3, 4, 1, 5]. Moreover, their materials and actuation methods are typically more affordable, safer, and simpler to manufacture or assemble compared to conventional rigid robots [6]. Additionally, soft robots have the ability to extend or retract through a process known as eversion – i.e., like a linear actuator, but with the advantage of occupying no space when completely retracted because the soft material can be folded. This unique degree of freedom (DoF) is directly inspired by the growth patterns of plants, leading to these robots often being referred to as vine robots [7, 8, 9] – or more generically soft growing robots. Several studies have demonstrated their effectiveness in manipulation tasks that require payload-handling capabilities [10, 11, 12, 13]. A major innovation was the introduction of continuum links and discrete joints, which can be achieved using layer jamming to dynamically adjust the robot’s stiffness at the joint locations to independently steer different links in different directions [14]. As illustrated in Fig. 1(a), this allows soft growing robots to be modeled with the classic kinematic chain used for their rigid counterparts, with the substantial difference that each link also grows and retracts.
Figure 1: (a) A soft growing manipulator with continuum links, variable stiffness, and variable discrete joints [14]. (b) Sketch of a three-dimensional soft growing robot reaching two targets in a maze (side view). The robot progressively grows, retracts, and steers to reach the next target. (c) A real prototype of a soft growing manipulator.
Soft robots are considerably less costly than their rigid counterparts, and this makes them accessible to the general public, including non-professional robot designers. Virtually anyone can create their own soft growing manipulator for their specific tasks [15, 16, 17]. However, specifying and integrating the components needed for a task-specific robot remains a complex challenge for human designers. With the capability of growing and retracting, the robot’s tip (namely the end effector or where the gripper/manipulative device is placed) can move anywhere along its kinematic chain. Consequently, their kinematic modeling and actuation pose significant difficulties. Ideally, a soft-robot designer would only need to define the functional goals and constraints of a given task, and an optimal design would be automatically generated [18].
Determining the design of a soft growing robot is a complex engineering problem that can be addressed using mathematical optimization [19]. A design defines the parameters and attributes describing a soft growing robot, ultimately guiding its manufacturing process. This includes aspects such as link lengths, joint positions and their range of motion (if applicable), materials used, and the overall shape or type of the robot. This problem has been addressed with several optimization methods: the interior point algorithm [20, 21, 22, 23, 24, 25], the Levenberg-Marquardt algorithm [20, 26], simplex methods [27, 28, 29, 30, 31, 32], the steepest descent method [33, 34], linear approximations [35], cyclic coordinate descent [36], Bayesian optimization [25], the method of moving asymptotes [37, 38], the Newton-Raphson method [39], greedy algorithms [40], hill climbing [41], simulated annealing [25], pattern search [42, 43], specific heuristics [44, 45], the parametric quadratic programming algorithm [46, 47], and the Broyden-Fletcher-Goldfarb-Shanno algorithm [48]. However, these traditional optimization methods often struggle with navigating highly complex, nonlinear, and multimodal design spaces, making it difficult to find global optima in engineering design problems. That is why engineering problems often rely on artificial intelligence, specifically the subfield of Evolutionary Computation (EC), a sub-field of Artificial Intelligence [49]. EC algorithms, which are natural-inspired metaheuristic optimizers, are widely used in mechanical and robotic design [50, 51, 52, 53, 54, 55]. Specifically for soft robotics, many studies have used EC algorithms to solve their design optimization problems: genetic algorithms for single [56, 18, 57, 58, 59, 60, 27, 61, 41] and multi-objective optimization [62, 63, 64, 65, 66, 67], particle swarm optimization [27, 68, 69, 70, 71], differential evolution [20], age-fitness-pareto optimization [72], the estimation of distribution algorithm [73], covariance matrix adaptation evolution strategy [61, 74, 75], and neuroevolution [65, 76]. These studies aim to optimize a soft robot’s design by finding the simplest possible configuration (i.e., the minimum number of sections) capable of performing a specific task or set of tasks [43]. The design optimization typically requires solving the robot’s kinematics, which, in the case of a soft manipulator, also involves optimizing its joint bending angles or curvature at each point in time [70].
Fig. 1(b) shows an example of a soft growing manipulator with discrete joints [14] specifically designed to solve a given manipulation task; whereas Fig. 1(c) shows a real picture of the manipulator. This specific problem has been addressed in the past by Exarchos et al. [17] in a two-dimensional scenario only – i.e., the manipulator acts in a planar environment. The same problem was addressed by our previous work [67], outperforming the state-of-the-art in terms of precision in reaching targets, smoothness of the configurations (i.e., non-wiggly or wavy movements to reach targets), and overall length of manipulators.
In this work, we extend our previous design optimizer for planar robots [67] into the third dimension to address more realistic scenarios with robots acting in a complex space – i.e., from two dimensions (2D) to three (3D). The design is modeled as a Multi-Objective Optimization Problem (MOOP) and addressed with EC algorithms. Specifically, we exploit the Rank Partitioning algorithm (introduced in the previous work [67]) to solve the MOOP without the need to retrieve the entire Pareto optimal front [77, 78]. This allows the optimizer to focus only on specific desired settings, thus minimizing the disadvantage of having trade-off designs.
Figure 2: Degrees of freedom (blue arrows) of the soft growing manipulator addressed in the current work, originally designed by Do et al. [14]. Each joint can rotate on x and y-axes, growing or retracting on the z-axis. Each joint has its own reference frame, whose z-axis is aligned in the direction of the previous link.
The contributions of this work are: (i) an extension of the innovative AI-based tool for robotic engineers presented in our previous work [67] by introducing the third dimension in the design; (ii) a mathematical formulation for this design optimization problem based on the kinematics proposed by Do et al. [14], which to the best of our knowledge has not been addressed in 3D before; (iii) the inclusion of additional constraints to the optimization problem that were overlooked or not needed in the 2D case; and (iv) a comparative robustness analysis of the optimizer and the Rank Partitioning algorithm when applied to four different EC algorithms.
The rest of the work is divided as follows: Section 2 describes the specific soft growing manipulator and the scenario we addressed; Section 3 offers a comprehensive overview of the optimization problem and a detailed walkthrough of the proposed solution utilizing EC algorithms; Section 4 presents an experiment to test the optimality of the method, particularly comparing how four different EC algorithms perform with Rank Partitioning; and Section 5 concludes the work illustrating possible future directions.
2 Soft Growing Manipulator for the Proposed Scenario
The soft growing manipulator examined in this study is based on the design introduced by Do et al. [14], which combines characteristics of both rigid and soft robots – Fig. 2 depicts a portion of the manipulator’s model to show its DoFs. The manipulator consists of links that can be individually adjusted to stiffen or stay flexible. When a link is not stiffened, it can be deliberately steered and maneuvered using cables to function as a revolute joint. When a link is stiffened, pulling its respective cable will not change its orientation; however, the following link (if not stiffeners) will be steered in the desired direction. Thanks to this technology, each link can steer independently by using only two cables for each direction (i.e., clockwise and anti-clockwise rotation around the x-axis, clockwise and anti-clockwise rotation around the y-axis). Lastly, growth and retraction (eversion) are made possible by the structure of the robot’s body, which is an inflatable beam composed of a thin-walled, compliant material that everts pneumatically. Starting in a fully folded state, pressurizing the interior with air causes the robot to unfold at its tip, pulling new material through its body [8, 79, 80].
Figure 3: Example of human-machine interaction with a three-dimensional soft growing robot. A soft manipulator mounted on the ceiling reaches down to help a human load/unload a dishwasher. The human’s role is to take dishes from the machine, whereas the robot’s role is to safely place them on the shelf.
This type of manipulator can be used in manipulation tasks [17] when equipped with a gripper at its tip (i.e., the end effector). This is usually done with a magnetic tip mount that moves as the robot grows, sliding on the body material, remaining anchored to the tip [81, 82, 11, 12]. In the specific scenario analyzed by this work, the robot is hanged on the ceiling and grows in favor of gravity to reach items and targets placed in the environment – much like a claw machine at the fair. Fig. 3 shows a possible real-case example in which the robot assists a human while unloading a dishwasher: the human selects the item from the dishwasher, hands it to the robot, and the robot places it in the designated spot on a shelf – a classic example of human-machine interaction. In this scenario, a soft robot is inherently safer than a rigid robot due to its flexibility and compliance [6]. Unlike rigid robots, which can exert high forces and have hard surfaces, soft robots are made of materials that can deform upon contact. This reduces the risk of injury if a collision with a human occurs. Additionally, the adaptable nature of soft robots allows them to handle objects of various shapes and fragilities with greater care, minimizing the chance of damaging items while also making the interaction more intuitive and safer for the user.
3 Optimization Problem and Proposed Solution
The problem requires us to find the best design for a soft growing robot manipulator to solve a specific task – i.e., the set of link lengths composing the body of the robot such that the robot can easily reach every target in the environment. Our optimizer will retrieve the design automatically. This problem can be defined in both 2D and 3D based on the effect of gravity on the system: a robot exploring a tunnel or a pipe system can be thought of as a 2D robot, but when the environment demands that the robot explore different levels (e.g., a multi-story building), it must also navigate in the third dimension. The 2D scenario was tackled in our previous work [67], and the current work extends it to the third dimension. Fig. 4 shows a 3D representation of the problem, where the robot is mounted on the environment’s ceiling and grows in favor of gravity [^1] In this work, gravity is oriented in the positive direction of the z-axis.. To simplify the problem, the model of our robot does not have gravity constraints on its orientation and can steer in any direction. Fig. 4(a) shows an example of a 3D task, in which the manipulator needs to reach three targets with a specified reaching orientation: this requires the robot to align its last link(s) to a specific straight line. Fig. 4(b) shows a possible solution to the problem, with three configurations (one for each target) sharing the same link lengths except for the last one.
(a) Task
(b) Solution
Figure 4: (a) Example of a task to be performed by the manipulator: three targets placed on different pillars with a given orientation. Starting from the home base $h$, the manipulator will grow such that its final link is aligned with each target. This task has three targets ($|t|=3$) and three obstacles ($|o|=3$) – i.e., the pillars where the targets are placed. (b) Example of solution $\psi$ provided by the optimizer. The solution $\psi$ comprises three configurations $\zeta_{1\rightarrow 3}$ sharing the same design $\delta$ (i.e., the lengths of links are shared among the three configurations). Each configuration can have a different number of joints, as the robot will employ only a sufficient number of links to reach a certain target. In this solution, the links whose endpoint corresponds to the tip of the robot are $\bar{n}_{1}=4$, $\bar{n}_{2}=5$, $\bar{n}_{3}=4$.
The robot begins its extension from its container, known as the home base. Based on the home base’s pose (position and orientation) and the set of targets to be reached, the optimizer determines an optimal robot design that allows it to reach all targets while satisfying hardware constraints (see Section 3.5). The design parameters include the number of joints and their placement along the robot: these parameters must be determined prior to the robot’s manufacturing and assembly – specifically, the lengths of the robot’s links. However, the number of joints or links that a soft growing robot can utilize during a task may change depending on the target it needs to reach. Additionally, the length of the final link may vary, as it depends on how much the robot has extended to reach a single target. Unlike rigid robots, where the design space is simply defined by the length of each link, the design space for a soft growing robot has unique characteristics due to its ability to grow and retract: (i) different targets may require varying numbers of links to be reached; (ii) interior links have a fixed length across different targets, while the length of the final link can change for each target but cannot exceed its maximum length (which equals to the length it would have if it were an interior link); (iii) a link can serve as either an interior or a final link for different targets; and (iv) the number of links required for a given set of targets is not predetermined and changes based on the task. Fig. 4(b) shows a single robot with fixed link lengths designed to reach three targets with a given pose: each robot configuration $\zeta_{i}$ (one for each target) has the same link lengths, except for the final one, as the specific target $t_{i}$ might not require additional extension, leaving the remainder of the robot folded within its body.
We model this scenario as an optimization problem: beginning with a random solution (random link lengths and joint angles for each configuration), we aim to minimize the distance between the robot’s tip and the targets’ poses using forward kinematics. Hence, the optimizer works as an inverse kinematics solver, ultimately determining the set of angles and link lengths (design) needed to reach each target.
3.1 Problem Formulation and Representation of the Decision Variables
The mathematical formulation of the problem defines the following terms:
  • 1.
    $n$, the maximum number of links;
  • 2.
    $\theta^{(x)}$, joint rotation around x-axis;
  • 3.
    $\theta^{(y)}$, joint rotation around y-axis; and
  • 4.
    $l$, link length.
Any start point of a link is defined as a node, which also corresponds to a set of two rotational joints (around x and y) and a linear joint (for growth and retraction). Only the last endpoint (i.e., the end effector or the node at the end of the last link) is not considered a joint set but simply the tip of the robot – i.e., not a node. Therefore, a robot will have at most $n$ nodes plus the end effector.
The user will set the bounds on these parameters based on the resources available for manufacturing – which might depend on budget, material availability, time, or the size of the machinery used for production. The number of nodes $n$ is defined in $\mathbb{N}^{+}$ and depends on how many links the manufacturer is prepared to produce. Angles on x and y are constrained within a range $\Delta_{\theta}\in\mathbb{R}$, which is determined by hardware limitations (e.g., each rotational joint cannot rotate more than $\pi/4$); this excludes the first joint at the base, where the robot extends orthogonally (i.e., either fixed at zero or with an unrestricted range if the home base is manually oriented for each target). Link lengths are constrained within a range $\Delta_{l}\in\mathbb{R}$ and are primarily influenced by manufacturing limitations (e.g., if the links are made from polyethylene plastic using a laser cutter, the maximum length is determined by the cutter’s size). Additionally, they must accommodate the gripping mechanism, meaning the link length must be at least as long as the gripper.
A specific task is defined with the following:
  • 1.
    $t$, a set of targets, each represented by their pose in the space (three coordinates for position and two angles for desired reaching orientation);
  • 2.
    $o$, a set of obstacles, each represented by their position in the space and other features – defined based on their geometry [^2] As described in Section 3.5, we represent all obstacles as cylinders oriented orthogonally with the floor, therefore the features of an obstacle include their radius, their height, and the coordinates of its origin as the center of the base.; and
  • 3.
    $h$, the pose of the home base (three coordinates for position and two angles for orientation) – although by default, we always place the home base to the origin of the reference frame, oriented toward the positive direction of the z-axis.
The result of the optimization problem consists of a design $\delta$, a solution $\psi$, and a set of $t$ configurations $\zeta_{i}$ (one for each target $i=1...|t|$):
  • 1.
    a single configuration $\zeta$ is defined as in Eqn. (1), and it is composed by the steering angles over x and y-axes at each joint and the length of each link;
    $\zeta=\begin{bmatrix}\theta^{(x)}_{1}&\theta^{(y)}_{1}&l_{1}\\ \vdots&\vdots&\vdots\\ \theta^{(x)}_{n}&\theta^{(y)}_{n}&l_{n}\\ \end{bmatrix}$ (1)
  • 2.
    multiple configurations define a solution [^3] A solution to an optimization problem is a well-defined set of values for the decision variables; this is also known as an individual in Evolutionary Computation. of the problem $\psi$, defined as in Eqn. (2) – note that each pair of rows represent steering angles on x and y, respectively, with the last row being reserved for shared link lengths (resulting in a $(2\cdot|t|+1)\times n$ matrix); and
[…]
  • 3.
    the desired design for the specific task is defined by the vector $\delta$ as in Eqn. (3), which contains the values of the link lengths of the robot (from the first to the last) – this is the last row of the matrix $\psi$ since all configurations share the same design.
    $\delta=\begin{bmatrix}l_{1}&\cdots&l_{n}\end{bmatrix}$ (3)
3.2 Search Space
A higher number of links enhances the robot’s dexterity and expands its workspace. Consequently, the optimization problem must handle a large search space $\omega$, which grows linearly with the number of targets $|t|$, as it needs to optimize a single set of links for all targets. The search space $\omega$ is defined as in Eqn. (4) when the order of targets is predetermined, or if the robot returns to its base after reaching each target (i.e., the order of targets is irrelevant).
$\omega=n^{\Delta_{l}}+2\cdot|t|\cdot n^{\Delta_{\theta}}$ (4)
When the order of targets is unknown, and the solution must also minimize the cost of transitioning between configurations, then the search space grows factorially, as every possible permutation of target sequences must be considered. In this work, we do not address the target order and focus solely on solving the basic optimization problem. The order can be considered after optimization with a proper motion planner like the one proposed by Altagiuri et al. [83] to determine the most efficient target-reaching sequence.
3.3 Fitness Evaluation from the Objective Functions
The optimization problem is composed of the following objectives:
  1. 1.
    $f_{1}(\psi,t,h,o)$: minimize the distance between the end effector (which can be located anywhere along the body due to eversion) and each target in the environment – solve the inverse kinematics;
  2. 2.
    $f_{2}(\psi,t,h,o)$: minimize the difference between the orientation of the robot’s last link and the desired reaching orientation for each target in the environment – solve the inverse kinematics; and
  3. 3.
    $f_{3}(\psi,t,h,o)$: minimize the robot’s components (i.e., its length, number of links), to promote quicker and more cost-effective production.
(b) Trade-off nodes/orientation
Figure 5: (a) Instead of directly aiming for the only target in the environment, this robot moves in random directions before completing the task, leading to a “wavy” configuration. (b) Trade-off problem between the number of nodes and reaching orientation. On the left, the robot reaches the target with the minimum possible number of nodes (it is assumed that all internal links are at their maximum length within the specified bounds), but it cannot reach the desired orientation. On the right, an additional node ensures that the robot achieves the correct orientation to reach the target.
In multi-objective optimization problems (MOOPs), some or all of the objectives may conflict with each other, meaning that optimizing one objective can adversely affect another. As a result, instead of finding a single solution, we typically obtain a set of trade-off solutions that represent the Pareto optimal front. Solutions on this front are non-dominated by any other solution, meaning no other solution is better across all objectives. However, as trade-off solutions, they may be optimal for one objective but not necessarily for others [77, 78]. In the context of this work, treating all objectives as equally important could result in trade-off solutions that are not desirable for our needs. For instance, while the minimum number of links might be a trivial single-link robot, it could be impractical if the robot cannot reach the target – i.e., a single-link robot cannot steer! Conversely, a robot with an excessive number of useless links (e.g., a “wavy” robot as shown in Fig. 5(a)) is clearly not the optimal solution we seek. Clearly, we are not interested in a robot that fails to reach the target, but we do need to balance the properties defining the robot (e.g., number of links) with its ability to achieve the desired orientation. Minimizing the number of links might compromise the ability to achieve the desired orientation, whereas an additional link can help meet the orientation requirement – as shown in Fig. 5(b). When dealing with multiple targets, achieving the exact desired orientation may be challenging, making trade-offs necessary. However, when feasible, it is preferable to use more links to better satisfy the orientation requirement – thus, $f_{2}$ takes precedence over $f_{3}$, and $f_{1}$ is given overall priority. Based on this, the optimization problem becomes to reach the target’s pose using the least amount of resources. Therefore, we should first verify that the configuration being examined can reach the target ($f_{1}$), and then proceed to minimize its components. It is important to note that although the robot must precisely reach the target, $f_{1}$ should not be treated as a constraint (i.e., solutions with a non-zero tip-to-target distance should not be deemed infeasible). If $f_{1}$ is treated as a constraint, the search algorithm may lose its heuristic effectiveness, leading to random convergence [^4] This claim is debatable, and it depends on the constraint-handling method employed: it would hold true for death or static penalty methods, but it could be valuable to explore results using dynamic methods or separation-based approaches [84], despite the higher risk of obtaining infeasible solutions only..
Using methods that retrieve Pareto optimal solutions, known as generating methods [85], may not be ideal for our situation unless we intentionally limit part of the objective space. Instead, we aim to integrate the three objectives into a single function, referred to as a fitness function. In addition to prioritizing the first objective, single-objective optimization methods generally have lower computational complexities than multi-objective methods.
To solve this design optimization problem, we employ the Rank Partitioning method, which was introduced in the optimizer for the 2D case [67]. Rank partitioning is a non-parametric survival strategy for Multi-Objective EC algorithms (or generically for population-based optimization methods) that ranks each individual based on a predefined order of the objectives. It sorts the population for the first objective (i.e., the one with the highest priority) and then creates partitions based on their objective rank; then, it sorts each partition for the second objective, and so on. Finally, it returns only one solution: the optimal based on the given order of priorities. It can be applied to multi or many-objective optimization (i.e., with more than three objectives [86]).
3.3.1 Solving the Inverse Kinematics with the Desired Orientation
The first two objectives are to ensure that the end effector of each configuration reaches the respective target ($f_{1}$) with the desired orientation ($f_{2}$). This is a well-known problem of inverse kinematics [87], which can be solved by minimizing the distance between the end effector and the target for each configuration. This is usually solved in task space (i.e., Cartesian coordinates in the environment) rather than in joint space (i.e., polar coordinates defining the values of each robotic joint). Therefore, since the configurations $\zeta$ in Eqn. (1) are defined in polar coordinates ($\zeta^{p}$), it is necessary to transform it into its corresponding configuration in Cartesian coordinates ($\zeta^{c}$) through forward kinematics as in Eqn. (5) – details are reported in Section 1.1 of the supplementary material. The configuration $\zeta^{c}$ is composed of $n+1$ points – i.e., all nodes plus the end effector.
$\zeta^{p}=\begin{bmatrix}\theta^{(x)}_{1}&\theta^{(y)}_{1}&l_{1}\\ \vdots&\vdots&\vdots\\ \theta^{(x)}_{n}&\theta^{(y)}_{n}&l_{n}\\ \end{bmatrix}\quad\xrightarrow[kinematics]{forward}\quad\zeta^{c}=\begin{bmatrix}x_{1}&y_{1}&z_{1}\\ \vdots&\vdots&\vdots\\ x_{n+1}&y_{n+1}&z_{n+1}\end{bmatrix}$ (5)
Figure 6: The phenotype of a robot generated with the genotype described in Eqn. (2), having $n=10$ links. In this case, only five are actually active to reach the target correctly ($\bar{n}=5$). Moreover, the last link is not fully extended, unlike the internal links that make up the rest of the body: this information is codified by extending the genotype as in Eqn. (6). The five links after the end effector (represented by dotted lines) do not need to be included in the design $\delta$.
(a) Inverse Kinematic Solver
(b) Phenotype Description
Figure 7: (a) When the inverse kinematic solver minimizes the distance node-target, there is no guarantee for the robot to have the last link aligned with the desired target orientation; however, when the solver minimizes the distance node-orientation, the robot will reach the desired orientation and can simply grow straight to the target after that. (b) The phenotype codified by a solution $\psi$ ignores every link after $\epsilon$ originally generated in the genotype and aims directly to the target for the remaining $\bar{n}-\epsilon$ links. The last link might not be fully everted, with a length smaller than the one originally generated in the genotype – indicated with the term $l_{\bar{n}}$.
The first issue to be addressed is the location of the end effector in a configuration: the population of any EC algorithm will be filled with random values for each solution $\psi$, meaning that the robotic configurations will explore the environment randomly to eventually converge with the evolutionary metaheuristic. Fig. 6 shows that the tip of the robot, or the end effector, can be placed anywhere on any link as the robot grows along its configuration. The link whose endpoint corresponds to the end effector is indicated with the index of $\bar{n}$. In case $n\neq\bar{n}$, in which $\bar{n}$ links are active and the rest $n-\bar{n}$ are folded within the robot, it is possible that the $\bar{n}$-th link is only partially everted – i.e., the target is reached before the link is fully grown. Based on the definition in Eqn. (2), multiple configurations share the same design because the robot can take different configurations with different angles, but the link lengths cannot change. However, when different targets require a different amount of growth, as shown in Fig. 4(b), it is necessary to modify the value of the last link for each configuration (i.e., $\bar{n}$) such that the link length only takes into account the active part of the link. In that case, the $\bar{n}$-th link length is not taken from the design $\delta$ (i.e., the last row of $\psi$ in Eqn. (2)), but is defined as the distance between the $\bar{n}$-th node and the target, and indicated with the term $l_{\bar{n}}$.
The second issue to be addressed is to guarantee that the configuration’s last link(s) lie(s) on the target’s desired orientation segment. The left-hand side of Fig. 7(a) shows that when minimizing the distance node $\bar{n}$ to target, the optimized configuration might end up not being aligned with the desired orientation. Rather than also minimizing the difference between the orientation of the robot’s last link and the desired one, it is more efficient to combine both objectives $f_{1}$ and $f_{2}$ into a single one, namely $f_{1-2}$. The right-hand side of Fig. 7(a) shows an alternative minimizing metric: reducing the distance between the end effector $\bar{n}+1$ to the target’s desired orientation segment. With this formulation, once the distance is minimized, the robotic configuration will intersect with the target’s orientation segment at the end effector $\bar{n}+1$. However, the end effector would still not be on the target, but somewhere on the segment: at this point, the robot can simply evert toward the target to eventually reach it – as shown in Fig. 7(b). This requires additional information to be stored in the genotype of a solution $\psi$. Firstly, the node selected to reach the target’s orientation segment is not the end effector but the closest node to the segment (at index $\epsilon$) among all the configuration nodes. Secondly, the rotation angles at the $\epsilon$-th joint (namely, $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$) must not follow the ones indicated in the configuration $\zeta$, but should be geometrically calculated such that the robot is oriented in the direction of the target (the procedure is reported in Sections 1.2 and 1.3 of the supplementary material). Fig. 7(b) shows the phenotype described by the genotype with the new extra information.
Considering these two issues, the additional information needs to be codified in the data structure/genotype of solutions $\psi$, whose elements define the decision variables of the optimization problem. The matrix $\psi$ then is defined as in Eqn. (6), where the left submatrix is the solution genotype $\psi$ defined in Eqn. (2), and the right submatrix is composed of the extra information for each configuration. Specifically, for each configuration, every pair of rows of the right submatrix contains:
  • 1.
    $\epsilon$, the index of the closest node to the target’s orientation segment;
  • 2.
    $\theta^{(x)}_{\epsilon}$, the angle on the x-axis for the $\epsilon$-th link of the robot to be aligned to the target’s orientation segment;
  • 3.
    $\theta^{(y)}_{\epsilon}$, the angle on the y-axis for the $\epsilon$-th link of the robot to be aligned to the target’s orientation segment;
  • 4.
    $\bar{n}$, the number of active nodes/links for each configuration (with $\bar{n}\leq n$) – namely the index of the last active node of a configuration; and
  • 5.
    $l_{\bar{n}}$, the length of the last link for each configuration, right before the end effector.
Note that, to keep a matrix shape, some elements are set to zero as they are unused: (i) the second row of each configuration, which defines rotations on the y-axis, only contains the last angle rotation $\theta^{(y)}_{\epsilon}$; and (ii) the last row, which defines the link lengths, are all fixed for a single design and do not need any extra information.
[…]
While the left submatrix is either randomly generated in the first generation of the EC algorithm or generated by variation when producing offspring, the right submatrix is filled during solution evaluation based on the geometrical requirements needed to reach each specific target. Also, it is important to note that the link lengths obtained in the design $\delta$ in Eqn. (3) serve as a reference only for the target farthest from the home base $h$, and that different configurations may use partial information from $\delta$ based on the number of active links.
The combined fitness $f_{1-2}$ is expressed as in Eqn. (7), where the distance between a node of the robot and the segment $s_{i}$ is calculated with the point-segment distance (introduced by Stroppa et al. [88] for 2D and extended for this work’s purposes in 3D as reported in Algorithm S.3 of the supplementary material). This computes the distance from the point to its projection onto the segment or to the nearest edge. This formulation also retrieves the index $\epsilon$ of the node that will intersect with the target’s orientation segment.
$f_{1-2}=\sum_{i=1}^{|t|}{\min_{\forall j\in n}{(\texttt{pointSegmentDistance}(\zeta^{c}_{i,j+1},s_{i})})}$ (7)
3.3.2 Optimizing the Kinematic Components of the Robot
This part is the same as the 2D case presented in the previous work [67], except for having two rotational angles for each node; however, we report it fully for a comprehensive description and reproducibility. The objective functions $f_{1}$ and $f_{2}$ (inverse kinematics) take precedence over $f_{3}$ (optimal components). After combining the first two into a single fitness function $f_{1-2}$, the next step is penalizing excessively long solutions with too many nodes. Ideally, we aim for a robot that gradually aligns with the target’s orientation segment, creating a smooth curve. The robot depicted in the right-hand side of Fig. 5(b) is a better solution compared to the one shown in Fig. 5(a): it not only positions all nodes closer to the orientation segment but also uses fewer nodes to reach the target.
Objective $f_{3}$ focuses on optimizing the robot’s components to minimize manufacturing time and resources and cannot be captured by a single property. Thus, we treat it as a macro objective composed of the following:
  • 1.
    $f_{3.1}$: minimize the number of links of each configuration to reduce redundant DoFs, as in Eqn. (8);
    $f_{3.1}=\sum_{i=1}^{|t|}\bar{n}_{i}$ (8)
  • 2.
    $f_{3.2}$: minimize the overall length of the robot to promote shorter configurations, as in Eqn. (9); and
    $f_{3.2}={\max_{\forall i\in|t|}{\left(\left(\sum_{j=1}^{\bar{n}_{i}-1}{l_{j}}\right)+l_{\bar{n}_{i}}\right)}}$ (9)
  • 3.
    $f_{3.3}$: the change in direction between links in all configurations, or undulation, to favor trajectories that are non-intersecting and not “wavy”, calculated as a percentage of the total length of all configurations, as shown in Eqn. (10) – note that the formula checks the change in the angle between the $j$-th and $(j+1)$-th link in both the $xz$ and $yz$-plane projections ($\Delta\theta^{(xz)}$ and $\Delta\theta^{(yz)}$, respectively).
    $f_{3.3}=\frac{\sum_{i=1}^{|t|}{\left(\frac{\forall k\in[xz,yz],\;\sum_{j=1}^{\epsilon_{i}-1}{\left[\operatorname{sign}{(\theta^{(k)}_{i,j})}\neq\operatorname{sign}{(\theta^{(k)}_{i,j+1})}\right]}}{2\cdot\epsilon_{i}}\right)}}{|t|}\cdot 100$ (10)
The priorities among these three sub-objectives can be adjusted to refine our preference-based method, potentially eliminating the need for weight parameters. Although it might appear that the number of links and the total length of the robot ($f_{3.1}$ and $f_{3.2}$, respectively) have a linear relationship (i.e., more links lead to a longer robot), this is not necessarily true. A robot with many small links can be shorter than one with fewer, very long links. We prefer robots with fewer links because the time required to manufacture a link is not affected by its length ($f_{3.1}$ takes precedence over $f_{3.2}$). On the other hand, the number of links and their orientations ($f_{3.1}$ and $f_{3.3}$, respectively) affect the robot’s shape differently. A robot with many links is more likely to be wavy, as illustrated in Fig. 5(a). Whether high-DoF non-wavy robots are preferable to low-DoF wavy ones may not be straightforward and could depend on the specific task (e.g., navigating a maze might require varying directions); however, we found that fewer links are preferable for manufacturing ($f_{3.1}$ takes precedence over $f_{3.3}$), but other manufacturers might have different priorities and choose to reverse this order. Finally, the undulation of a robot and its length ($f_{3.3}$ and $f_{3.2}$, respectively) are strongly correlated. A robot that directly targets the destination will generally be shorter than one that moves through the environment before reaching the target, as demonstrated in Fig. 5(a). Therefore, we consider it more important to construct non-wavy robots, which will ultimately lead to shorter configurations ($f_{3.3}$ takes precedence over $f_{3.2}$).
Furthermore, we found that to enhance the quality of the solutions obtained, the objective $f_{3.1}$ should be broken down into two sub-objectives:
  • 1.
    $f_{3.1a}$: minimize the number of links to reach the target’s orientation segment, as in Eqn. (11); and
    $f_{3.1a}=\sum_{i=1}^{|t|}(\epsilon_{i}-1)$ (11)
  • 2.
    $f_{3.1b}$: positioned along the target’s orientation segment, as in Eqn. (12).
    $f_{3.1b}=\sum_{i=1}^{|t|}(\bar{n}_{i}-\left(\epsilon_{i}-1\right))$ (12)
By making this division, we specifically direct how the configuration should be arranged rather than just minimizing the total number of links across different configurations. This involves prioritizing the reduction of links required to align with the target’s orientation segment. Once this segment is reached, we then focus on minimizing the number of links needed to reach the target itself. Note that the final step can be effectively achieved by manually maximizing the lengths of these links. Although the optimization algorithm should ideally learn this heuristic, practical implementation might require some manual adjustments due to the extensive search space.
3.4 Separation of the Objectives with Rank Partitioning
1
input : $N$, population size.
input : $F$, a matrix $N\times 9$ having the following information for each individual: fitness for each objective (integer values first followed by real values), fitness after partitioning ($rank$), reference to the individual in the population ($ref$). A single row of $F$ is: [(1)[(1)[ ( 1 ): $F_{1-2\mathbb{N}}$, $(2)$: $F_{3.1a}$, $(3)$: $F_{3.3}$, $(4)$: $F_{3.1b}$, $(5)$: $F_{3.2\mathbb{N}}$, $(6)$: $F_{1-2\mathbb{R}}$, $(7)$: $F_{3.2\mathbb{R}}$, $(8)$: $rank$, $(9)$: $ref]_{i}$ $\forall i\in[1,N]$.
input : $binSize$, distance between partitions for each real-coded objective (i.e., one for $F_{1-2}$ and one for $F_{3.2}$).
output : The updated matrix $F$ with the value of $rank$ after partitioning.
2
3begin
4       for $i\leftarrow 1$ to $N$ do
5             $F_{i,1}\leftarrow F_{i,6}-\textnormal{{mod}}(F_{i,6},binSize_{1-2})$
6             $F_{i,5}\leftarrow F_{i,7}-\textnormal{{mod}}(F_{i,7},binSize_{3.2})$
7
8
9      sortrows($f,>$) on column 1-5 for all rows
10       $d\leftarrow 0_{N\times 5}$
11       for $i\leftarrow 2$ to $N$ do
12             for $j\leftarrow 1$ to $5$ do
13                   $d_{i,j}\leftarrow\textnormal{{abs}}(F_{i,j}-F_{i-1,j})$
14
15
16
17      $start\leftarrow 1$
18       for $i\leftarrow 1$ to $N$ do
19             for $j\leftarrow 1$ to $5$ do
20                   $partLimit\leftarrow partLimit+d_{i,j}$
21
22            if $partLimit>0$ then
23                   $end\leftarrow i$
24                   sortrows($f,>$) on column 6-7 from row start to end-1
25                   $start\leftarrow end$
26
27
28
29      for $i\leftarrow 1$ to $N$ do
30             $F_{i,8}\leftarrow i$
31
32      return $F$
33
Algorithm 1 Rank Partitioning [67]
Based on the observations about objective priorities, we employ a strategy that prioritizes solutions with fewer, well-structured components, but only if they are sufficiently close to the target’s orientation segment. The Rank Partitioning algorithm was introduced in the previous work by Stroppa [67] and is used to retrieve the best solution of a MOOP given a predetermined order of priorities among objectives, without the burden of compromising the optimality of the highest prioritized objectives and choosing weights for each of them (i.e., unlike weighted sum methods [78, 89]).
Rank partitioning divides the population into $k$ partitions (bins) with fixed intervals (bin sizes) according to their kinematic fitness $f_{1-2}$. Individuals with better fitness are assigned to the first segments, while those with worse fitness are placed in the last segments – in this specific context, where all objectives will be minimized, the best fitness is the one with the lowest value. This is repeated for each objective: the partitions made for the first objective will be divided into more partitions based on the second objective, and so on, until all the objectives are partitioned. The key idea is that individuals within the same partition are considered equally optimal in terms of the partitioned objective so that the best among them can be selected based on the following objectives. Just like in the 2D scenario [67], we applied Rank Partitioning in the following order for each partition:
  1. 1.
    distance from the target/inverse kinematics ($f_{1-2}$);
  2. 2.
    number of links to reach the targets’ orientation segment ($f_{3.1a}$);
  3. 3.
    percentage of undulation among configurations ($f_{3.3}$);
  4. 4.
    number of links on the targets’ orientation segment ($f_{3.1b}$); and
  5. 5.
    length of the robot/final design ($f_{3.2}$).
This approach works only when ties happen within partitions, leading to nested sub-partitions that follow the prioritized order of objectives. This creates a challenge when handling objectives with continuous domains, as we deal with ranges or bins of fixed intervals rather than individual values (similar to a histogram). To address the issue, Rank Partitioning duplicates all continuous objectives (i.e., $f_{1-2}$ and $f_{3.2}$) by adding a further objective composed by the lower-bound discrete value of their bin. Discrete objectives can be used with no modification (i.e., $f_{3.1}\in[1,n\cdot|t|]$ and $f_{3.3}\in[0,100]$). In this manner, the partitioning operation can be resolved using a straightforward sorting algorithm that accounts for ties. When a tie occurs in the objective with the lowest priority, individuals are ranked based on their original continuous values of continuous-domain objectives (i.e., $f_{1-2}$ and $f_{3.2}$), which should be preserved rather than substituted by their discrete bin values. Finally, Rank Partitioning updates the fitness value by assigning a rank to each individual in the population, which increases linearly according to their position within the partitions (i.e., sequentially numbering each individual after the partitioning process from 1 to $N$, with $N$ being the population size).
Algorithm 1 shows Rank Partitioning on the soft growing robot problem, which runs in $\mathcal{O}(N\texttt{log}N)$, where $N$ represents the population size. The input $F$ is a “fitness” matrix $N\times 9$ containing, for each individual, (i) its kinematic fitness $f_{1-2}$, (ii) the four sub-objectives describing the components $f_{3}$, (iii) the original values of the continuous objectives that were evaluated by their lower-bound bin value, (iv) the updated rank value, calculated by the procedure, will be utilized by the selection operators for the reproduction and survival of individuals, and (v) a reference to the individual in the population [^5] The reference is used to link fitness in the matrix and the corresponding individual in the population since only the fitness matrix will be sorted to save computational time. This is simply a pointer to the respective individual in the population.. Lines 2-4 convert continuous objectives (denoted by the subscript $\mathbb{R}$) into discrete values (denoted by subscripts $\mathbb{N}$) based on the specified bin size [^6] Note that the subscript $\mathbb{N}$ is just a placeholder since these values are not limited to natural numbers only as their bin size can be smaller than 1.. Line 5 sorts the rows of the fitness matrix according to the five discrete objectives in the order provided. Lines 6-9 determine partition boundaries by computing differences between consecutive elements in each sorted column of the fitness matrix: if two consecutive elements in the same column are equal, they belong to the same partition. Summing these differences for each row (lines 12-13) helps the algorithm identify partition boundaries. Lines 11-17 re-sort each row by the original continuous values, but only within partitions, ensuring the order from line 5 remains intact. Finally, lines 18-20 assign the updated rank to each individual.
In multi-objective optimization, this approach acts as a restriction within the objective space, focusing only on the segment of the Pareto optimal front with the lowest values for one objective: the first partition represents the optimal solution for $f_{1-2}$, allowing us to select the individual that minimizes $f_{3}$ as the best fit. The key benefits of this method are: (i) it enforces priority on a specific objective (e.g., inverse kinematics in our case), (ii) it eliminates the need to select weight values as in typical preference-based methods, and (iii) it provides flexibility for future research to incorporate additional sub-objectives that may be relevant but were not considered in this study (i.e., for tackling many-objective optimization problems).
3.5 Constraints
The problem’s constraints are similar to the 2D case presented in the previous work [67], with the respective additions of one more dimension for the rotational angles and one more constraint on obstacle collision that was not previously – and erroneously – considered. Every configuration in the solutions is subject to eight kinematic constraints in addition to the bounds on decision variables (i.e., $\Delta_{\theta}$ and $\Delta_{l}$, which are parameters the user sets prior to manufacturing).
Since the final links (i.e., from the $\epsilon$-th to the $\bar{n}$-th node) are automatically determined to reach the target as described in Section 3.3.1, the orientation angles $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$ between the last active link and the target’s orientation segment depends on the kinematic chain produced by the genetic operators. As a result, it may exceed the bounds $\Delta_{\theta}$, as illustrated in Fig. 8(a). We defined four inequality constraints, as shown in Eqn. (13), to mark solutions as infeasible when configurations have $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$ beyond these limits. Note that the bound $\Delta_{\theta}$ can also differ between axes, but for the sake of simplicity, we will report it as the same value.
Figure 8: Some examples of infeasible solutions. In (a), excessive steering when reaching the target; in (b), a very small link at the $\epsilon$-th link right after steering that cannot accommodate a gripper device; in (c), not reaching the target with the desired orientation; in (d), an obstacle collision; and in (e) a collision non detected in the configuration but as result of the motion planner, which requires a link to be everted before steering.
$\theta^{(x)}_{\epsilon}\geq{\Delta_{\theta}}^{(L)}\wedge\theta^{(x)}_{\epsilon}\leq{\Delta_{\theta}}^{(U)}$ (13)
$\theta^{(y)}_{\epsilon}\geq{\Delta_{\theta}}^{(L)}\wedge\theta^{(y)}_{\epsilon}\leq{\Delta_{\theta}}^{(U)}$
Similarly, the length of each configuration is determined by the distance between the $i$-th target and the $\bar{n}_{i}$-th node, which may result in configurations where the length of the final link, $l_{\bar{n}}$, falls below its lower bound limit. In practice, this becomes an issue only when the final link needs to adjust its direction to align with the target’s orientation segment. Considering that these manipulators are likely equipped with a gripping mechanism at the tip (e.g., the tip mount proposed by Jeong et al. [81] or the mechanism to prevent buckling proposed by Coad et al. [82]), it is essential to account for the necessary space to position the gripper properly on the target’s orientation segment for accurate grasping. We formulated the bounds ${\Delta_{l}}^{(L)}$ to account for this constraint. As a result, any solutions featuring configurations where a single link on the target’s orientation segment ($f_{3.1b}=1$, Eqn. (12)) is insufficiently long are considered infeasible, as illustrated in Fig. 8(b). This leads to the definition of a fifth inequality constraint, as expressed in Eqn. (14).
$\text{if}\;f_{3.1b}=1\text{,}\quad l_{\bar{n}}\geq{\Delta_{l}}^{(L)}$ (14)
The $\epsilon$-th node, identified as the nearest joint to the target’s orientation segment, is determined using the point-segment distance method (details are reported in Section 1.2 of the supplementary material). Since this process may designate one of the segment’s endpoint as the nearest point to $\epsilon$ – rather than $\epsilon$’s projection on the segment – if that endpoint corresponds to the target, the resulting solution would resemble the configuration shown in Fig. 8(c). Specifically, we refer to the case at line 19 of Algorithm S.3 presented in the supplementary material. To avoid this issue, we established an additional inequality constraint as defined in Eqn. (15). This constraint checks whether the angle width between the end of each configuration $\zeta_{i}$ and the target’s orientation segment $s_{i}$ achieves or exceeds $\pi/2$. Specifically, the target’s orientation segment $s_{i}$ (mentioned in Eqn. (7) as well) has two endpoints: $s_{i,1}$ is the coordinates of its respective target $t_{i,1\rightarrow 3}$, and $s_{i,2}$ is any point on the straight line where $s_{i}$ lies. The angle between these two segments must be smaller than $\pi/2$ for the configuration to be feasible; otherwise, it means that the robot overshot and is trying to align with the target rather than its orientation segment.
$\cos^{-1}\left(\frac{(\zeta^{c}_{\epsilon_{i}}-s_{i,1})\cdot(s_{i,2}-s_{i,1})}{\|\zeta^{c}_{\epsilon_{i}}-s_{i,1}\|\cdot\|s_{i,2}-s_{i,1}\|}\right)<\frac{\pi}{2},\text{ where }i\in[1,...,|t|]$ (15)
Lastly, we impose two constraints related to obstacle collisions. If the environment contains elements that are not targets, they must be treated as obstacles. Any solution that results in a collision with these obstacles is marked as infeasible, as illustrated in Fig. 8(d). In our optimization problem, obstacle collisions are treated as an equality constraint, formulated as shown in Eqn. (16).
$\texttt{intersections}(\psi,o)=0$ (16)
The function in Eqn. (16) calculates the number of intersections between each obstacle in the set $o$ and every link of each configuration in $\psi$. The full procedure is reported in Section 1.4 of the supplementary material, and it operates with a complexity of $\mathcal{O}(\overline{n}\cdot|o|\cdot|t|)$. To simplify the process and reduce computational overhead, we model every obstacle as a cylinder due to the ease of detecting intersections between a segment and a cylinder during a collision [90]. More complex obstacles can be modeled by combining multiple cylinders: a pillar is represented by a single cylinder; a wall by a line of cylinders; and a hole in a wall by stacking cylinders in a bottom layer, a top layer, and two taller cylinders at the sides, forming a hole in the middle. During problem setup, users can select a radius slightly larger than that of the obstacle’s circumscribed cylinder to prevent contact between the robot and the obstacle. Note that avoiding obstacles might also be considered as another objective function: one can (i) minimize obstacle collisions or (ii) maximize the distance between any robot link and obstacles. However, the first option would consider a solution with obstacle collision as feasible, which deceives our purposes as we do not want collisions to happen at all. On the other hand, the second option might not work if a target is surrounded by obstacles (e.g., the target is in an open box). Therefore, considering obstacle collisions as a constraint is the best option for our optimization problem.
The last constraint regards a possible obstacle collision when actuating the manipulator – i.e., the static configuration $\zeta_{i}$ does not collide with obstacles, but the actuation to move to $\zeta_{i}$ from any other configuration $\zeta_{j}$ will cause a collision. This is because of how the manipulator is actuated: a link must grow of at least a minimum length (i.e., the lower bound $\Delta^{(L)}_{l}$) before is enabled to steer. Similarly to the constraint expressed in Eqn. (14) and Fig. 8(b), this prevents the robot from steering while the joint is within the tip mount, avoiding possible damage. Fig. 8(e) shows an example: if the configuration requires the robot to grow of a length equal to $\Delta^{(L)}_{l}$ within the obstacle, then we have a collision – note that the figure represents a projection in 2D of a spherical cone. Note that a motion planner could actually find a feasible alternative path without colliding with the obstacle (i.e., by bending one of the previous joints); however, checking this during optimization would require running the motion planner [83], which would require too much in terms of computational resources. Therefore, we simply mark a configuration like this one as infeasible. This constraint is summarized in Eqn. (17), which imposes the number of collisions while steering to be equal to zero. More details are provided in Section 1.5 of the supplementary material.
$\texttt{cannotSteer}(\psi,o)=0$ (17)
4 Experimental Results and Discussion
This section illustrates the experiments to (i) assess whether the optimizer can retrieve optimal solutions on tasks with large dimensionality and many obstacles; (ii) determine which EC algorithm is the best fit for solving our design problem among four different options; and (iii) evaluate the robustness of the Rank Partitioning algorithm when included in the loop of different EC algorithms.
input : The optimization problem $op$ (in our case, described by $t$, $h$, $o$, $n$, $\Delta_{\theta}$, $\Delta_{l}$) including a set of constraints $c$
input : The EA parameter such as number of individuals $N$ and number of generations $G$
output : The most fitting individual $best$ and its fitness $f^{best}$
1
2begin
3
4      $pop\leftarrow\textnormal{{randomInitialization}}(N)$
5       $F^{pop}\leftarrow\textnormal{{evaluation}}(op,c,pop)$
6       $F^{pop}\leftarrow\textnormal{{rankPartitioning}}(F^{pop})$
7       for $i\leftarrow 1$ to $G$ do
8             $off\leftarrow\textnormal{{variation}}(pop,F^{pop})$
9             $F^{off}\leftarrow\textnormal{{evaluation}}(op,c,off)$
10
11            if apply elitism then
12                   $F^{pop}\cup F^{off}\leftarrow\textnormal{{rankPartitioning}}(F^{pop}\cup F^{off})$
13                   $pop,F^{pop}\leftarrow\textnormal{{survival}}(pop,off,F^{pop}\cup F^{off})$
14
15             else
16                   $F^{off}\leftarrow\textnormal{{rankPartitioning}}(F^{off})$
17                   $pop\leftarrow off$
18                   $F^{pop}\leftarrow F^{off}$
19
20            $best,f^{best}\leftarrow\textnormal{{rankPartitioning}}(f^{best}\cup F^{pop}_{1})$
21
22      return $best$, $f^{best}$
23
24
Algorithm 2 EC framework with Rank Partitioning
4.1 Optimization Methods
We tested the design optimizer on four different EC algorithms: Genetic Algorithm (GA) [91], Particle Swarm Optimization (PSO) [92], Differential Evolution (DE) [93], and Big Bang-Big Crunch algorithm (BBBC) [94]. Each of these EC algorithm has its own parameters, which are described in the following subsections; however, some of the parameters are shared as they all comply with the basic framework of EC algorithms, reported in Algorithm 2: an initial population $pop$ is initialized with $N$ random individuals (line 2), evaluated returning the fitness of each individual $F^{pop}$ penalized in case of constraint violation (line 3), ranked based on the Rank Partitioning algorithm (line 4, runs Algorithm 1), and then it undergoes the generational loop for $G$ iterations (line 5-14) in which variation operators generate a new population of offspring $off$ (line 6), evaluate them returning their fitness $F^{off}$ with penalty in case of constraint violation (line 7), and select which individuals among $pop$ and $off$ will survive and be part of the next generation based on the survival scheme (lines 8-14) – specifically, when elitism is applied, the Rank Partitioning algorithm is executed on both populations (line 9), otherwise only on the offspring (line 12). Note that, with Rank Partitioning, the best individual to be saved and returned by the method is not the individual with minimum/maximum fitness value, but it is the one having rank $=1$ after executing Rank Partitioning between the previous best and the best of the current generation (line 15) – regardless on whether the method is elitist or not. According to preliminary evaluations [^7] For better readability, we do not report any explicit detail of this preliminary study except for the p-values. However, the data can be provided at readers’ request., we set only GA to be elitist (with elitism being significantly better than non-elitist, $p<0.01$), whereas the other EC algorithms were all set to be non-elitist – PSO and DE were found to be significantly better when elitism is not applied ($p<0.01$ and $p=0.0179$, respectively), and no significant difference was observed with BBBC ($p=0.3793$). Table 1 reports the parameters and settings shared among all EC algorithms, with the values we used in the study.
Table 1: Experimental conditions and parameters (shared among EC algorithms)
Parameter Value
Population Size ($N$) $500$
Max Number of Generations ($G$) $500$
Constraint Handling Static Penalty [95, 84] $R=100$
Rank Partitioning Bin Size for $f_{1-2}=0.5$ for $f_{3.2}=5.0$
The population size $N$ was chosen based on the results of the previous 2D work [67], whereas the maximum number of generations $G$ was set to a high number to ensure the convergence of the algorithms – i.e., the value of the best solution retrieved by the optimizer stopped changing before the run terminated.
For constraint handling, we used the Static Penalty method [95, 84]. With this method, infeasible solutions are penalized with a fixed penalty factor $R$. We empirically set the value of $R$ to $100$ for every constraint, a value that allows the optimizer to retrieve solutions without fluctuating between feasible and infeasible at each iteration.
Finally, we chose Rank Partitioning’s bin sizes for the continuous-to-discrete objectives based on the results of the previous 2D work [67]. Specifically, for $f_{1-2}$ (inverse kinematics), a greater bin size corresponds to a significant decrease in the number of links; however, for some tasks where flexibility is required, a smaller bin size can be preferred. This allows the robot to maneuver comfortably between obstacles. For $f_{3.2}$ (total robot length), we observed the average total link lengths from the twenty pre-runs we used to estimate the convergence generation ($200\pm 78$) and compared them to the ones in the 2D work ($216\pm 63$). Since these robot lengths are similar, we selected the same bin size value.
4.1.1 Genetic Algorithm (GA)
GAs are designed for optimization in continuous or discrete domains by simulating the process of natural evolution [91]. In GAs, candidate solutions, known as individuals, are typically represented as encoded vectors or chromosomes. The algorithm iteratively evolves the population through genetic operations: selection, crossover, and mutation. During selection, individuals with higher fitness are chosen to propagate their traits to the next generation. Crossover combines pairs of selected individuals to create offspring, allowing for the exploration of new regions in the search space. Mutation introduces small, random changes to individual genes, helping to maintain diversity and avoid premature convergence. By iteratively applying these operations, GAs balance exploration and exploitation of the search space, guiding the population toward optimal or near-optimal solutions over successive generations. This flexibility makes GAs particularly effective for complex, multimodal optimization problems.
Table 2: Experimental conditions and parameters (GA)
Parameter Value
Selection Type Binary Tournament [96]
Crossover Type blx-$\alpha$ [97] $\alpha=0.500,0.419,0.381$
Crossover Probability ($p_{c}$) 1.0
Mutation Type Random [98]
Mutation Probability ($p_{m}$) 0.2, 0.4, 0.6, Dynamic $1.0\rightarrow 0.0$
Survival Type Elitist ($\mu+\lambda$ schema) [99]
Table 2 reports the parameter settings we used for GA. We used the binary tournament for selecting individuals undergoing variation [96], the blx-$\alpha$ for crossover [97], random mutation [98], and both elitist and non-elitist methods for survival. Specifically, we tested three values for $\alpha$ in the crossover, which dictates how far the offspring are generated from the parents – these values were selected based on a comparative study on self-adaptation among crossover methods [100] to determine the range of probability distribution between parents. Since the effects of mutation cannot be predicted on the specific problem (i.e., what happens with different degrees of exploration), we tested four different probabilities of mutation, including a strategy that dynamically and linearly decreases it from 1 to 0 with the generation counter (i.e., full exploration at early generations, full exploitation toward the end). The elitist survival method employs the $\mu+\lambda$ schema [99], which (i) merges the old population and the offspring, (ii) sorts them by fitness (in our case, with Rank Partitioning), and (iii) selects only the $N$ best individuals.
4.1.2 Particle Swarm Optimization (PSO)
PSO is designed for optimization in continuous domains by representing each candidate solution, or particle, as a position in a multi-dimensional space [92]. Unlike other algorithms that rely on crossover or mutation, PSO uses the concept of velocity to iteratively update the positions of particles. Each particle adjusts its velocity based on both its own previous best position and the best-known positions of its neighbors or the entire swarm. This dynamic adjustment allows particles to explore the search space more effectively. The search directions in PSO are influenced by both local and global information, promoting a balance between exploration and exploitation. This process enables particles to converge toward optimal or near-optimal solutions while avoiding getting trapped in local optima, making PSO particularly effective for solving complex, nonlinear optimization problems.
Table 3: Experimental conditions and parameters (PSO)
Parameter Value
Inertia Weight ($\omega$) $0.6,0.8,1.0$ [101]
Cognitive Constant ($c_{1}$) $0.5,2.5,5.0$ [102, 103]
Social Constant ($c_{2}$) $0.5,2.5,5.0$ [102, 103]
Table 3 reports the parameter settings we used for PSO. The inertia weight $\omega$ is between [0.5, 1.0] after the random $\omega$ factor formula [101]. As for $c_{1}$ and $c_{2}$, Ratnaweera et al. [103] reported the best range for those constants being [0.5, 2.5], with 5.0 being the best parameter suggested by Zhou et al. [102]
4.1.3 Differential Evolution (DE)
DE is specifically designed for optimization in continuous domains by representing each potential solution, or individual, as a multi-dimensional vector [93]. Perturbations are introduced into the population through vector differences, allowing DE to explore the search space efficiently. A key distinction from GAs lies in its mutation strategy: instead of relying on predefined probability distributions or crossover rates, DE generates mutations based on the differences between randomly selected vectors from the current population. This dynamic and adaptive approach allows the search directions to be guided by the relative positions of individuals, promoting diverse exploration and reducing the chances of premature convergence to local optima. Additionally, DE’s mechanism of mutation and recombination maintains the balance between exploration and exploitation, which is crucial for solving complex, nonlinear optimization problems.
Table 4: Experimental conditions and parameters (DE)
Parameter Value
Variant Type rand/1, best/1, rand/2, best/2, current-to-best/1, current-to-rand/1  [104]
Scaling Factor ($F$) $0.50,0.75,1.00$  [93]
Table 4 reports the parameter settings we used for DE. We chose the six most widely used DE variants [104]. Regarding the scaling factor $F$, we chose values between 0.50 and 1.00 since 0.5 is a good initial choice and values below 0.4 or above 1.00 are usually not effective [93].
4.1.4 Big Bang-Big Crunch Algorithm (BBBC)
BBBC is designed for optimization in continuous domains by simulating the evolutionary process of the universe’s expansion (Big Bang) and contraction (Big Crunch) [94]. In the Big Bang phase, candidate solutions, represented as points in a multi-dimensional space, are randomly scattered to explore the search space. During the Big Crunch phase, these points are drawn toward a central point (usually the best solution found so far) to exploit promising regions. This contraction of solutions guides the search directions, allowing the algorithm to refine and intensify the exploration around the most optimal areas. BBBC dynamically balances exploration and exploitation by iterating between these two phases, enhancing its ability to escape local optima and converge toward global optima, making it effective for solving complex, nonlinear optimization problems.
BBBC does not feature any parameter, so the only settings we used are the shared ones reported in Table 1.
(a) Task 1
(b) Task 2
(c) Task 3
Figure 9: Representation of each task in both isometric (left) and top view (right), along with their respective optimal solutions retrieved by the optimizer.
4.2 Tasks
We selected three tasks with progressively increasing difficulty for which we could estimate that an optimal solution exists. In all tasks, the robot base is hung on the ceiling and reaches the targets from the top – specifically, at 150 cm from the ground [^8] The unit of measurement is irrelevant, as the problem can be scaled based on the size of the workspace; however, the dimensions we tried are based on realistic measurements in our environment, so we report them in cm.. The tasks are:
  • 1.
    Task 1: two targets placed on two pillars and three additional pillars as obstacles, a simple task in which the obstacles do not hinder the reachability of the targets (shown in Fig. 9(a));
  • 2.
    Task 2: four targets placed on the ground (i.e., a desk) and four pillars as obstacles, a medium task in which the obstacles are in the path of the targets (shown in Fig. 9(b)); and
  • 3.
    Task 3: eight targets placed on eight pillars and eight additional pillars as obstacles, a hard task in which the robot must find openings within this “wall” to reach many targets (shown in Fig. 9(c)).
We executed each task with a maximum number of nodes $n=20$, bounds on joint rotation set to $\Delta_{\theta}=\pi/4$ on each direction (a limited value emulating hardware constraints), and bounds on link lengths set to $\Delta_{l}\in[25,70]$ to fit the workspace. We repeated each task fifty times for each condition to test all the combinations of parameters discussed in Section 4.1 and summarized in the following Section 4.3.
4.3 Parameter Combinations
The experiments present the following parameter combinations:
  • 1.
    3 tasks (see Section 4.2);
  • 2.
    4 optimization algorithms (see Section 4.1), of which:
    • (a)
      12 combinations for GA: 3 values for blx-$\alpha$ and 4 mutation probabilities (see Section 4.1.1 and Table 2);
    • (b)
      27 combinations for PSO: 3 values for inertia weight, 3 cognitive constants, and 3 social constants (see Section 4.1.2 and Table 3);
    • (c)
      18 combinations for DE: 6 variants, and 3 scaling factors (see Section 4.1.3 and Table 4); and
    • (d)
      1 combination for BBBC (see Section 4.1.4);
    summing up to the 58 parameter combinations reported in Table S.1 of the supplementary material; and
  • 3.
    50 runs for each combination.
This sums up to $58\cdot 50=2900$ runs for each task and $2900\cdot 3=8700$ runs overall.
4.4 Evaluation Metric
Since we are dealing with a five-objective optimization problem (three objectives in which one of them is further divided into two, as described in Section 3.3), we used the rank from Rank Partitioning as the evaluation metric. The Rank Partitioning algorithm, described in Section 3.4, sorts the retrieved solutions based on their objective values according to their predefined priorities; then, it assigns a rank to each of them that linearly increments in ascending order from best to worst – i.e., the best solution gets rank $=1$, the second best gets rank $=2$, and so on.
Therefore, we evaluated the results by collecting the $2900$ retrieved solutions for each task and assigned them a rank with Rank Partitioning – i.e., from $1$ to $2900$.
4.5 Results
Fig. 10 reports the rank results for each task and each EC algorithm, including median, interquartile range with outliers, and max/min across all runs. Separate results for each objective are reported in the supplementary material in Figure S.2, whereas a complete rank comparison among each of the 58 parameter combinations is depicted in Figure S.3. Table 5 reports the numerical values of the best solution retrieved by each algorithm in each task, where the asterisk indicates the solution at rank 1 depicted in Fig. 9.
Figure 10: Ranking according to Rank Partitioning, over three tasks and each EC algorithm. Optimality is indicated by low-rank values (i.e., minimize rank). The same plot, divided among all parameter combinations among each EC algorithm, is shown in Fig. S.3 of the supplementary material.
Table 5: Objective values of the best designs retrieved by each EA. The overall best (according to Rank Partitioning and depicted in Figure 9) is indicated with a .
IK ($f_{1-2}$) LtS ($f_{3.1a}$) UND ($f_{3.3}$) LoS ($f_{3.1b}$) LEN ($f_{3.2}$) Rank Constraints
Task 1 GA $0.00$ cm $6$ $50\%$ $2$ $198.68$ cm $32/2900$ feasible
PSO* $0.25$ cm $6$ $25\%$ $4$ $197.57$ cm $1/2900$ feasible
DE $0.43$ cm $6$ $25\%$ $4$ $196.97$ cm $8/2900$ feasible
BBBC $0.37$ cm $6$ $25\%$ $4$ $197.90$ cm $5/2900$ feasible
Task 2 GA* $0.26$ cm $12$ $31\%$ $6$ $224.87$ cm $1/2900$ feasible
PSO $0.50$ cm $12$ $38\%$ $4$ $226.49$ cm $8/2900$ feasible
DE $0.42$ cm $12$ $56\%$ $5$ $234.35$ cm $264/2900$ feasible
BBBC $0.42$ cm $12$ $44\%$ $8$ $221.81$ cm $89/2900$ feasible
Task 3 GA* $0.31$ cm $27$ $52\%$ $21$ $301.99$ cm $1/2900$ feasible
PSO $2.98$ cm $28$ $57\%$ $22$ $302.47$ cm $436/2900$ feasible
DE $12.96$ cm $26$ $54\%$ $19$ $295.20$ cm $612/2900$ feasible
BBBC $7.99$ cm $28$ $45\%$ $22$ $291.71$ cm $565/2900$ feasible
We adopted Friedman’s test with the Bonferroni-Dunn procedure [105] to test the robustness of each EC algorithm on the 58 parameter combinations for a total of 1653 pairwise comparisons (results are reported in the supplementary material under Tables S.2–S.5). The p-values show the result of the Bonferroni-Dunn procedure (ordered from Task 1 to 3 from top to bottom for each pairwise comparison); “Y” indicates that there is a significant difference (at the significance level $\alpha=0.05$) between the corresponding pairwise settings, whereas “N” indicates no significant difference.
4.6 Discussion
The optimal designs shown in Fig. 9 (whose values are reported in Table 5) suggest that the optimizer is able to solve this complex engineering problem with no visible challenge. While Task 1 is easy and does not require specific tuning or a specific EC algorithm to be solved, the other tasks are more complex. Yet, all the designs can precisely reach the targets (precision is reported under $f_{1-2}$ in Table 5), show smooth and non-wavy configurations, with a small number of links (i.e., minimizing material waste). Furthermore, all the solutions are feasible, meaning that each algorithm was able to satisfy the constraints imposed in Section 3.5. The most surprising outcome, however, is that the configurations in a single design are almost mirrored. This was unexpected as EC algorithms are stochastic algorithms, and each configuration in the same design is not affected by the other in terms of joint actuation. We purposely chose mirrored tasks (i.e., all obstacles and targets are placed in the same fashion on the left-hand side as well as on the right-hand side) such that a mathematically optimal solution would retrieve a perfect mirrored solution. The top-view plots in Fig. 9 show that our designs are very close to being mirrored, which is evidence of the optimality of our method: this was particularly unexpected in Task 3, and we could claim that no artificial fix-ups are required to smooth these configurations (i.e., manually remove the right-hand side configurations and replace them with a mirrored copy of the left-hand side ones).
In terms of EC algorithms, the numerical results show that GA is consistently more efficient in retrieving a better solution. This is visible from the plots in Fig. 10: in all three tasks, the solutions from GA occupy the first ranks according to Rank Partitioning. This is even more visible on Task 3, which is the hardest of the three: while the other EC algorithms occasionally retrieve low-rank solutions in the first two tasks (e.g., the best solution for Task 1 was retrieved with PSO as shown in Table 5), GA retrieved all the solutions from rank $1$ to $435$ in Task 3 – meaning that the best $15\%$ solutions were all retrieved by GA – whereas the worst algorithm for this type of problem is DE.
Specifically, according to the Friedman’s test and Bonferroni-Dunn procedure (Tables S.2–S.5 of the supplementary material), there are no significant differences among all the GA settings, in all the tasks. This indicates that no specific parameter setting of Table 2 is more suitable than another, and designers can simply select a GA with any of these values. Furthermore, we found no statistical difference, in all tasks, between the following (with reference to the IDs in Table S.1 of the supplementary material):
  • 1.
    all GA settings and PSO 4 ($\omega=0.6$, $c_{1}=2.5$, $c_{2}=0.5$);
  • 2.
    all GA settings and PSO 13 ($\omega=0.8$, $c_{1}=2.5$, $c_{2}=0.5$);
  • 3.
    four out of twelve GA settings and DE 6 (best/1, $F=1.0$); and
  • 4.
    three out of twelve GA settings and DE 15 (current-to-best/1, $F=1.0$).
This indicates that GA, PSO 4, and PSO 13 are the best algorithms for our soft growing robot design. Regarding BBBC, it was never found to be non-significantly different than any GA settings in all three tasks (for Task 2, GA was always significantly better). However, it was found to be non-significantly different than PSO 4, PSO 13, DE 6, and DE 15 (the settings that were non-significantly different than GA). By looking at the plots in Fig. 10, we could claim that BBBC is, on average, more consistent than PSO and DE; although specific settings for the latter two often outperform it. This is interesting, considering that BBBC is a less-known metaheuristic, which yet could compete with big names in the state-of-the-art.
In conclusion, we found that our optimizer can retrieve high-performance 3D designs in terms of precision in reaching targets, resource consumption, and actuation, especially when the core EC algorithm is GA.
4.7 Considerations on Problem Dimensionality
The tasks described in Section 4.2 are optimization problems with high dimensionality in the variable search space. This can be quantified from the fixed value of the maximum link number $n=20$ and the number of targets $|t|$ in the task. Each link accounts for two decision variables per joint (i.e., rotation on $x$ and $y$), which is multiplied for each desired target (i.e., each configuration of the robot aims at one specific target in the environment. The link lengths are shared among all configurations, so they add a constant term. This results in a dimensionality of $2\cdot n\cdot|t|+n$. We can then evaluate the following dimensionalities:
  • 1.
    Task 1 (Fig. 9(a)): $2\cdot 20\cdot 2+20=100$ decision variables;
  • 2.
    Task 2 (Fig. 9(b)): $2\cdot 20\cdot 4+20=180$ decision variables; and
  • 3.
    Task 3 (Fig. 9(c)): $2\cdot 20\cdot 8+20=340$ decision variables.
In the previous 2D work [67], the algorithm was tested only up to 140 decision variables. The tasks analyzed in the current work are inherently more complex due to the third dimension of the environment; yet, as shown by the results reported in Section 4.5, the optimizer can tackle high dimensionality problems while still achieving (sub)optimal designs – especially when the core EC algorithm is GA. It is reasonable to claim that the limit on dimensionality depends on the computational cost of the specific task.
5 Conclusion
In this work, we presented an extension of our innovative tool to optimize the design of soft growing robots for solving specific tasks. While the previous work was limited only to planar manipulators (i.e., in 2D), the current work completes the tool addressing more realistic 3D scenarios. We used the Rank Partitioning algorithm, introduced in our previous work, to solve a multi-objective optimization problem without compromising among trade-offs. Since this algorithm can theoretically work on every population-based optimizer, we tested it on four different Evolutionary Computation techniques: Genetic Algorithm, Particle Swarm Optimization, Differential Evolution, and Big Bang-Big Crunch algorithm. Results show that the proposed optimizer retains the outstanding performances of its previous 2D version, especially when based on a Genetic Algorithm.
In future works, we will use the optimizer on a real specific task, manufacture the soft robot according to the retrieved design, and use our motion planner [83] to demonstrate the robot’s capabilities as a manipulator. However, this poses two new challenges that could not be addressed in this work: (i) a proper control module is required to address not only kinematic actuation but also dynamic components, and (ii) the amount of material loss in the kinematic chain when links bend, which is a real-case problem that was not modeled in this work. Additionally, we plan to include a further objective in the optimization problem: minimize the path between configurations such that the robot visits all targets with the optimized order – also known as Traveling Salesman Problem [106]. This objective cannot be handled directly in the optimization due to its extreme computational complexity; however, it is useless to calculate the optimal order for solutions far away from the optimum (i.e., high-rank solutions according to Rank Partitioning). Therefore, this objective could be evaluated only for a subset of the individuals in the population having low rank, reducing the amount of computation required [107].
Acknowledgments
This work is funded by TUBİTAK within the scope of the 2232-B International Fellowship for Early Stage Researchers Program number 121C145.
The authors would like to thank: Prof. Allison M. Okamura for inspiring the project, Kadir Kaan Atalay and Kemal Erdem Yenin for writing the code of BBBC, Kerem Bicen for writing the code of DE, Ayodele Oyejide for providing pictures of a real soft growing robot, and Emre Ozel for allowing us to use 121 computers in the Kadir Has University’s Computer Labs for parallel computation.
Supplementary Information
A document containing the supplementary material can be downloaded at https://www.mediafire.com/file/s4r3yqm489cccg6/3D+SGR+supplementary+material.pdf/file.
A graphical presentation on the Rank Partitioning algorithm, including theory and hands calculation, is available at https://www.mediafire.com/file/6jlp1yr4p8yjjul/22._Rank_Partitioning_for_MOEAs.pdf/file.
Readers interested in manufacturing these soft growing robots can find more information at https://www.vinerobots.org/.

### v2

@@
 Soft robotics is a relatively new field of intelligent mechanical systems that aims at overcoming some of the biggest limitations of their rigid counterparts [1].
-Rigid robots (i) may struggle to interact with fragile objects or delicate environments, such as human tissues, glassware, or sensitive components in archaeological sites; (ii) have intricate designs and their materials often result in high costs; and (iii) due to their strength and stiffness, the safety of human operators can be a concern in a human-machine interaction scenario [2].
+Rigid robots (i) may struggle to interact with delicate objects or environments, such as human tissues, glass objects, or sensitive components in archaeological sites; (ii) have intricate designs and their materials often result in high costs; and (iii) due to their strength and stiffness, the safety of human operators can be a concern in a human-machine interaction scenario [2].
 On the other hand, soft robots are inspired by living organisms, with designs exploiting the natural compliance of their materials to facilitate smooth and intuitive physical interactions with the environment [3, 4, 1, 5].
-Moreover, their materials and actuation methods are typically more affordable, safer, and simpler to manufacture or assemble compared to conventional rigid robots [6].
+Moreover, their materials and actuation methods are typically safer, cheaper, and simpler to manufacture or assemble compared to conventional rigid robots [6].
 Additionally, soft robots have the ability to extend or retract through a process known as eversion – i.e., like a linear actuator, but with the advantage of occupying no space when completely retracted because the soft material can be folded.
-This unique degree of freedom (DoF) is directly inspired by the growth patterns of plants, leading to these robots often being referred to as vine robots [7, 8, 9] – or more generically soft growing robots.
-Several studies have demonstrated their effectiveness in manipulation tasks that require payload-handling capabilities [10, 11, 12, 13].
-A major innovation was the introduction of continuum links and discrete joints, which can be achieved using layer jamming to dynamically adjust the robot’s stiffness at the joint locations to independently steer different links in different directions [14].
+This unique degree of freedom (DoF) is inspired by the growth patterns of plants, leading to these robots often being referred to as vine robots [7, 8, 9] – or more generically soft growing robots.
+Several studies have demonstrated their effectiveness in manipulation tasks that require handling of payloads [10, 11, 12, 13].
+A major innovation was introduced by Do et al.
+[14] with continuum links and discrete joints: thanks to dynamic layer jamming, the robot can independently stiffen the different parts of its body, resulting in a joint.
 As illustrated in Fig.
 1(a), this allows soft growing robots to be modeled with the classic kinematic chain used for their rigid counterparts, with the substantial difference that each link also grows and retracts.
-Figure 1: (a) A soft growing manipulator with continuum links, variable stiffness, and variable discrete joints [14].
-(b) Sketch of a three-dimensional soft growing robot reaching two targets in a maze (side view).
-The robot progressively grows, retracts, and steers to reach the next target.
-(c) A real prototype of a soft growing manipulator.
-Soft robots are considerably less costly than their rigid counterparts, and this makes them accessible to the general public, including non-professional robot designers.
-Virtually anyone can create their own soft growing manipulator for their specific tasks [15, 16, 17].
+Figure 1: (a) A soft growing manipulator with continuum links and variable discrete joints [14].
+(b) Sketch of a three-dimensional soft growing robot reaching two targets in an environment with obstacles (side view).
+(c) A picture of a physical manipulator.
+Soft robots are considerably less costly than their rigid counterparts, and this makes them accessible to everyone, including amateur roboticists.
+Every robotic enthusiast could create their soft growing manipulator for one or more well-defined specific tasks [15, 16, 17].
 However, specifying and integrating the components needed for a task-specific robot remains a complex challenge for human designers.
@@
 Consequently, their kinematic modeling and actuation pose significant difficulties.
-Ideally, a soft-robot designer would only need to define the functional goals and constraints of a given task, and an optimal design would be automatically generated [18].
-Determining the design of a soft growing robot is a complex engineering problem that can be addressed using mathematical optimization [19].
+Preferably, a designer only needs to define the task itself, including any hardware or environmental constraint, and use an optimizer to retrieve optimal designs [18].
+Determining a soft growing robot’s design is a complex engineering problem that mathematical optimization can easily address [19].
 A design defines the parameters and attributes describing a soft growing robot, ultimately guiding its manufacturing process.
@@
 These studies aim to optimize a soft robot’s design by finding the simplest possible configuration (i.e., the minimum number of sections) capable of performing a specific task or set of tasks [43].
-The design optimization typically requires solving the robot’s kinematics, which, in the case of a soft manipulator, also involves optimizing its joint bending angles or curvature at each point in time [70].
+The design optimization typically requires solving the robot’s kinematics, which, in the case of a soft manipulator, also involves finding the correct curvature along the body [70].
 Fig.
@@
 This allows the optimizer to focus only on specific desired settings, thus minimizing the disadvantage of having trade-off designs.
+The contributions of this work are: (i) an extension of the innovative AI-based tool for robotic engineers presented in our previous work [67] by introducing the third dimension in the design; (ii) a mathematical formulation for this design optimization problem based on the kinematics proposed by Do et al.
+[14], which to the best of our knowledge has not been addressed in 3D before; (iii) the inclusion of additional constraints to the optimization problem that were overlooked or not needed in the 2D case; and (iv) a comparative robustness analysis of the optimizer and the Rank Partitioning algorithm when applied to four different EC algorithms.
+The rest of the work is divided as follows: Section 2 describes the specific soft growing manipulator and the scenario we addressed; Section 3 offers a comprehensive overview of the optimization problem and a detailed walkthrough of the proposed solution utilizing EC algorithms; Section 4 presents an experiment to test the optimality of the method, particularly comparing how four different EC algorithms perform with Rank Partitioning; and Section 5 concludes the work illustrating possible future directions.
 Figure 2: Degrees of freedom (blue arrows) of the soft growing manipulator addressed in the current work, originally designed by Do et al.
@@
 Each joint has its own reference frame, whose z-axis is aligned in the direction of the previous link.
-The contributions of this work are: (i) an extension of the innovative AI-based tool for robotic engineers presented in our previous work [67] by introducing the third dimension in the design; (ii) a mathematical formulation for this design optimization problem based on the kinematics proposed by Do et al.
-[14], which to the best of our knowledge has not been addressed in 3D before; (iii) the inclusion of additional constraints to the optimization problem that were overlooked or not needed in the 2D case; and (iv) a comparative robustness analysis of the optimizer and the Rank Partitioning algorithm when applied to four different EC algorithms.
-The rest of the work is divided as follows: Section 2 describes the specific soft growing manipulator and the scenario we addressed; Section 3 offers a comprehensive overview of the optimization problem and a detailed walkthrough of the proposed solution utilizing EC algorithms; Section 4 presents an experiment to test the optimality of the method, particularly comparing how four different EC algorithms perform with Rank Partitioning; and Section 5 concludes the work illustrating possible future directions.
 2 Soft Growing Manipulator for the Proposed Scenario
-The soft growing manipulator examined in this study is based on the design introduced by Do et al.
+The soft growing manipulator examined in this study is based on the design proposed by Do et al.
 [14], which combines characteristics of both rigid and soft robots – Fig.
@@
 Lastly, growth and retraction (eversion) are made possible by the structure of the robot’s body, which is an inflatable beam composed of a thin-walled, compliant material that everts pneumatically.
-Starting in a fully folded state, pressurizing the interior with air causes the robot to unfold at its tip, pulling new material through its body [8, 79, 80].
+Starting in a fully folded state, pressurizing the interior with air unfolding the tip, pulling new material through its body [8, 79, 80].
 Figure 3: Example of human-machine interaction with a three-dimensional soft growing robot.
@@
 Additionally, the adaptable nature of soft robots allows them to handle objects of various shapes and fragilities with greater care, minimizing the chance of damaging items while also making the interaction more intuitive and safer for the user.
-3 Optimization Problem and Proposed Solution
+(a) Task
+(b) Solution
+Figure 4: (a) Task example: three targets placed on different pillars with a desired orientation.
+The robot grows from its base $h$ such that its final link is aligned with each target.
+This task has three targets ($|t|=3$) and three obstacles ($|o|=3$) – i.e., the pillars where the targets are placed.
+(b) Optimizer’s output example: the solution $\psi$ comprises three configurations $\zeta_{1\rightarrow 3}$ that share a unique design $\delta$ (i.e., the three configurations $\zeta$ share the same link lengths).
+Because the robot may use only a limited amount of links to reach a specific target, depending on its distance to it, each configuration can have a different number of employed links.
+In this example, the number of links whose endpoint corresponds to the robot’s tip are $\bar{n}_{1}=4$, $\bar{n}_{2}=5$, $\bar{n}_{3}=4$.
+3 Optimization Problem
+This section describes the optimization problem, its formalization, and how the objectives were engineered to formulate a proper fitness evaluation.
+Since the details are similar to the 2D case, with the exception of adding a further dimension for the rotational joints, we briefly summarize this description and invite readers to find more details in our previous work [67].
+However, unlike our previous work, more information has been added regarding the specifics of the procedures, which can be found in Section 1.1 of the Supplementary Material.
 The problem requires us to find the best design for a soft growing robot manipulator to solve a specific task – i.e., the set of link lengths composing the body of the robot such that the robot can easily reach every target in the environment.
 Our optimizer will retrieve the design automatically.
-This problem can be defined in both 2D and 3D based on the effect of gravity on the system: a robot exploring a tunnel or a pipe system can be thought of as a 2D robot, but when the environment demands that the robot explore different levels (e.g., a multi-story building), it must also navigate in the third dimension.
+We defined this problem in both 2D and 3D based on the effect of gravity on the system: a robot navigating a tunnel or a pipe system can be thought of as 2D; but when the environment demands navigating different levels or layers, it requires 3D exploration.
 The 2D scenario was tackled in our previous work [67], and the current work extends it to the third dimension.
 Fig.
-4 shows a 3D representation of the problem, where the robot is mounted on the environment’s ceiling and grows in favor of gravity [^1] In this work, gravity is oriented in the positive direction of the z-axis..
-To simplify the problem, the model of our robot does not have gravity constraints on its orientation and can steer in any direction.
+4 shows a 3D representation of the problem, where the robot is mounted on the environment’s ceiling and grows in favor of gravity – oriented in the positive direction of the z-axis.
+To simplify the problem, we did not consider gravity in the model.
 Fig.
-4(a) shows an example of a 3D task, in which the manipulator needs to reach three targets with a specified reaching orientation: this requires the robot to align its last link(s) to a specific straight line.
-Fig.
-4(b) shows a possible solution to the problem, with three configurations (one for each target) sharing the same link lengths except for the last one.
-(a) Task
-(b) Solution
-Figure 4: (a) Example of a task to be performed by the manipulator: three targets placed on different pillars with a given orientation.
-Starting from the home base $h$, the manipulator will grow such that its final link is aligned with each target.
-This task has three targets ($|t|=3$) and three obstacles ($|o|=3$) – i.e., the pillars where the targets are placed.
-(b) Example of solution $\psi$ provided by the optimizer.
-The solution $\psi$ comprises three configurations $\zeta_{1\rightarrow 3}$ sharing the same design $\delta$ (i.e., the lengths of links are shared among the three configurations).
-Each configuration can have a different number of joints, as the robot will employ only a sufficient number of links to reach a certain target.
-In this solution, the links whose endpoint corresponds to the tip of the robot are $\bar{n}_{1}=4$, $\bar{n}_{2}=5$, $\bar{n}_{3}=4$.
-The robot begins its extension from its container, known as the home base.
-Based on the home base’s pose (position and orientation) and the set of targets to be reached, the optimizer determines an optimal robot design that allows it to reach all targets while satisfying hardware constraints (see Section 3.5).
-The design parameters include the number of joints and their placement along the robot: these parameters must be determined prior to the robot’s manufacturing and assembly – specifically, the lengths of the robot’s links.
-However, the number of joints or links that a soft growing robot can utilize during a task may change depending on the target it needs to reach.
-Additionally, the length of the final link may vary, as it depends on how much the robot has extended to reach a single target.
-Unlike rigid robots, where the design space is simply defined by the length of each link, the design space for a soft growing robot has unique characteristics due to its ability to grow and retract: (i) different targets may require varying numbers of links to be reached; (ii) interior links have a fixed length across different targets, while the length of the final link can change for each target but cannot exceed its maximum length (which equals to the length it would have if it were an interior link); (iii) a link can serve as either an interior or a final link for different targets; and (iv) the number of links required for a given set of targets is not predetermined and changes based on the task.
-Fig.
-4(b) shows a single robot with fixed link lengths designed to reach three targets with a given pose: each robot configuration $\zeta_{i}$ (one for each target) has the same link lengths, except for the final one, as the specific target $t_{i}$ might not require additional extension, leaving the remainder of the robot folded within its body.
-We model this scenario as an optimization problem: beginning with a random solution (random link lengths and joint angles for each configuration), we aim to minimize the distance between the robot’s tip and the targets’ poses using forward kinematics.
-Hence, the optimizer works as an inverse kinematics solver, ultimately determining the set of angles and link lengths (design) needed to reach each target.
-3.1 Problem Formulation and Representation of the Decision Variables
-The mathematical formulation of the problem defines the following terms:
+4(a) shows an example of a 3D task in which the manipulator needs to reach three targets according to a desired reaching orientation.
+The robot should then align its last link – or more than one – so that it lies on the requested straight line.
+One example of design is shown in Fig.
+4(b), with three configurations (one for each target) sharing the same link lengths except for the last one.
+The robot begins its extension from its container, known as the home base $h$.
+Based on the home base’s pose (position and orientation) and the environment, the optimizer determines the optimal robot design so that the manipulator can reach every target without violating any constraint (see Section 3.3).
+The joint number and placement define the parameters of the design, which must be known prior to the robot’s manufacturing and assembly – specifically, the design $\delta$ is a set of link lengths defining a single manipulator.
+This results in a single optimal solution $\psi$ having multiple configurations $\zeta$ for each target (i.e., a different manipulator’s orientation reaching each target) sharing a unique link-length design $\delta$.
+Configurations can differ from each other due to eversion: they may use a different link number because not all the links may actually be employed – i.e., they are folded within the robot’s body.
+Additionally, it is also not guaranteed that the last link will be completely everted, depending on how far the last employed joint is from the specific target.
+We model this scenario as an optimization problem: beginning with a random solution (random link lengths and joint angles for each target/configuration) generated through forward kinematics, it minimizes the distance from the end effector and each of the targets’ poses.
+The design (link lengths and, consequently, link angles) are retrieved as an inverse kinematic solver.
+Table 1: Nomenclature of the terms defined in the Optimization Problem
+Symbol Description
+$\delta$ design
+$\psi$ solution
+$\zeta$ configuration
+$n$ maximum number of links/nodes
+$\theta^{(x)}$ joint rotation around x-axis
+$\theta^{(y)}$ joint rotation around y-axis
+$l$ length of link
+$t$ set of targets
+$o$ set of obstacles
+$h$ home base pose
+$\epsilon$ index of the closest node to target’s orientation segment
+$\theta^{(x)}_{\epsilon}$ joint rotation around x-axis at the $\epsilon$-link toward target
+$\theta^{(y)}_{\epsilon}$ joint rotation around y-axis at the $\epsilon$-link toward target
+$\bar{n}$ number of active/folded links
+$l_{\bar{n}}$ length of the last active/folded link
+$f_{1-2}$ objective function: inverse kinematics
+$f_{3.1a}$ objective function: number of links to segment
+$f_{3.1b}$ objective function: number of links on segment
+$f_{3.2}$ objective function: overall robot length
+$f_{3.3}$ objective function: undulation
+The optimization problem is defined with the terminology summarized in Table 1.
+3.1 Problem Formulation
+We begin the formulation with the following:
   • 1.
-    $n$, the maximum number of links;
+    $n\in\mathbb{N}^{+}$, maximum link number, corresponding to the number of links the user is prepared to produce;
   • 2.
-    $\theta^{(x)}$, joint rotation around x-axis;
+    $\theta^{(x)}$ and $\theta^{(y)}\in\mathbb{R}$, rotation of joint around x and y-axes, respectively, and constrained within a domain $\Delta_{\theta}$, which is determined by how much the robot can bend without breaking or buckling – note that we decided to exclude the joint corresponding to the home base (i.e., the first) due to hardware observations so that the robot will grow from the base in the direction of its orientation and steer only from the second joint on; and
   • 3.
-    $\theta^{(y)}$, joint rotation around y-axis; and
-  • 4.
-    $l$, link length.
+    $l\in\mathbb{R}$, length of a link, constrained within a domain $\Delta_{l}$, and similar to the maximum link number $n$, it is influenced by production limitations (i.e., the overall size of the material sheet used for the body, the size of the cutting tool, etc.).
+Additionally, they have to be longer than the gripping mechanism to ensure stability on the end effector – i.e., imagine the gripper that moves along the body as the robot everts, pushed by the end effector [81]; this mechanism has a specific length, which should be smaller than any link length because if a joint happens to be inside the gripper, that joint cannot bend.
 Any start point of a link is defined as a node, which also corresponds to a set of two rotational joints (around x and y) and a linear joint (for growth and retraction).
-Only the last endpoint (i.e., the end effector or the node at the end of the last link) is not considered a joint set but simply the tip of the robot – i.e., not a node.
-Therefore, a robot will have at most $n$ nodes plus the end effector.
-The user will set the bounds on these parameters based on the resources available for manufacturing – which might depend on budget, material availability, time, or the size of the machinery used for production.
-The number of nodes $n$ is defined in $\mathbb{N}^{+}$ and depends on how many links the manufacturer is prepared to produce.
-Angles on x and y are constrained within a range $\Delta_{\theta}\in\mathbb{R}$, which is determined by hardware limitations (e.g., each rotational joint cannot rotate more than $\pi/4$); this excludes the first joint at the base, where the robot extends orthogonally (i.e., either fixed at zero or with an unrestricted range if the home base is manually oriented for each target).
-Link lengths are constrained within a range $\Delta_{l}\in\mathbb{R}$ and are primarily influenced by manufacturing limitations (e.g., if the links are made from polyethylene plastic using a laser cutter, the maximum length is determined by the cutter’s size).
-Additionally, they must accommodate the gripping mechanism, meaning the link length must be at least as long as the gripper.
-A specific task is defined with the following:
+Only the last endpoint (i.e., the end effector) is not considered a joint set but simply the robot’s tip – i.e., not a node.
+A design will have, at most, $n$ nodes plus the end effector.
+The problem formulation also includes the task, which is defined as:
   • 1.
-    $t$, a set of targets, each represented by their pose in the space (three coordinates for position and two angles for desired reaching orientation);
+    $t$, target set, defined as an array of poses (i.e., for each target, three coordinates for position and two angles for desired reaching orientation);
   • 2.
-    $o$, a set of obstacles, each represented by their position in the space and other features – defined based on their geometry [^2] As described in Section 3.5, we represent all obstacles as cylinders oriented orthogonally with the floor, therefore the features of an obstacle include their radius, their height, and the coordinates of its origin as the center of the base.; and
+    $o$, obstacle set, defined as an array of poses and geometrical properties [^1] In this work, obstacles are modeled as cylinders oriented orthogonally with the floor, described by a radius, height, and the coordinates of its origin as the center of the base – see in Section 3.3.; and
   • 3.
-    $h$, the pose of the home base (three coordinates for position and two angles for orientation) – although by default, we always place the home base to the origin of the reference frame, oriented toward the positive direction of the z-axis.
-The result of the optimization problem consists of a design $\delta$, a solution $\psi$, and a set of $t$ configurations $\zeta_{i}$ (one for each target $i=1...|t|$):
+    $h$, home base pose (three coordinates for position and two angles for orientation) – although by default, we always place the home base to the origin of the reference frame, oriented toward the positive direction of the z-axis.
+The output of the optimization consists of the following terms:
   • 1.
-    a single configuration $\zeta$ is defined as in Eqn. (1), and it is composed by the steering angles over x and y-axes at each joint and the length of each link;
+    configurations $\zeta_{i}$ (one for each target $i=1...|t|$), containing a column vector for the steering angles over the x-axis, y-axis, and the length of each link as in Eqn. (1);
     $\zeta=\begin{bmatrix}\theta^{(x)}_{1}&\theta^{(y)}_{1}&l_{1}\\ \vdots&\vdots&\vdots\\ \theta^{(x)}_{n}&\theta^{(y)}_{n}&l_{n}\\ \end{bmatrix}$ (1)
   • 2.
-    multiple configurations define a solution [^3] A solution to an optimization problem is a well-defined set of values for the decision variables; this is also known as an individual in Evolutionary Computation.
-of the problem $\psi$, defined as in Eqn.
+    a solution $\psi$ contains $|t|$ configurations, which provides the values of the decision variables and a description of the individuals’ genotype, defined as in Eqn.
 (2) – note that each pair of rows represent steering angles on x and y, respectively, with the last row being reserved for shared link lengths (resulting in a $(2\cdot|t|+1)\times n$ matrix); and
@@
   • 3.
-    the desired design for the specific task is defined by the vector $\delta$ as in Eqn.
-(3), which contains the values of the link lengths of the robot (from the first to the last) – this is the last row of the matrix $\psi$ since all configurations share the same design.
+    the final design $\delta$ defined as in Eqn. (3), containing link lengths as they appear in $\psi$ – i.e., one design for multiple configurations in the solution.
     $\delta=\begin{bmatrix}l_{1}&\cdots&l_{n}\end{bmatrix}$ (3)
-3.2 Search Space
-A higher number of links enhances the robot’s dexterity and expands its workspace.
-Consequently, the optimization problem must handle a large search space $\omega$, which grows linearly with the number of targets $|t|$, as it needs to optimize a single set of links for all targets.
-The search space $\omega$ is defined as in Eqn.
-(4) when the order of targets is predetermined, or if the robot returns to its base after reaching each target (i.e., the order of targets is irrelevant).
-$\omega=n^{\Delta_{l}}+2\cdot|t|\cdot n^{\Delta_{\theta}}$ (4)
-When the order of targets is unknown, and the solution must also minimize the cost of transitioning between configurations, then the search space grows factorially, as every possible permutation of target sequences must be considered.
-In this work, we do not address the target order and focus solely on solving the basic optimization problem.
-The order can be considered after optimization with a proper motion planner like the one proposed by Altagiuri et al.
-[83] to determine the most efficient target-reaching sequence.
-3.3 Fitness Evaluation from the Objective Functions
-The optimization problem is composed of the following objectives:
-  1. 1.
-    $f_{1}(\psi,t,h,o)$: minimize the distance between the end effector (which can be located anywhere along the body due to eversion) and each target in the environment – solve the inverse kinematics;
-  2. 2.
-    $f_{2}(\psi,t,h,o)$: minimize the difference between the orientation of the robot’s last link and the desired reaching orientation for each target in the environment – solve the inverse kinematics; and
-  3. 3.
-    $f_{3}(\psi,t,h,o)$: minimize the robot’s components (i.e., its length, number of links), to promote quicker and more cost-effective production.
-(b) Trade-off nodes/orientation
-Figure 5: (a) Instead of directly aiming for the only target in the environment, this robot moves in random directions before completing the task, leading to a “wavy” configuration.
-(b) Trade-off problem between the number of nodes and reaching orientation.
-On the left, the robot reaches the target with the minimum possible number of nodes (it is assumed that all internal links are at their maximum length within the specified bounds), but it cannot reach the desired orientation.
-On the right, an additional node ensures that the robot achieves the correct orientation to reach the target.
-In multi-objective optimization problems (MOOPs), some or all of the objectives may conflict with each other, meaning that optimizing one objective can adversely affect another.
-As a result, instead of finding a single solution, we typically obtain a set of trade-off solutions that represent the Pareto optimal front.
-Solutions on this front are non-dominated by any other solution, meaning no other solution is better across all objectives.
-However, as trade-off solutions, they may be optimal for one objective but not necessarily for others [77, 78].
-In the context of this work, treating all objectives as equally important could result in trade-off solutions that are not desirable for our needs.
-For instance, while the minimum number of links might be a trivial single-link robot, it could be impractical if the robot cannot reach the target – i.e., a single-link robot cannot steer!
-Conversely, a robot with an excessive number of useless links (e.g., a “wavy” robot as shown in Fig.
-5(a)) is clearly not the optimal solution we seek.
-Clearly, we are not interested in a robot that fails to reach the target, but we do need to balance the properties defining the robot (e.g., number of links) with its ability to achieve the desired orientation.
-Minimizing the number of links might compromise the ability to achieve the desired orientation, whereas an additional link can help meet the orientation requirement – as shown in Fig.
-5(b).
-When dealing with multiple targets, achieving the exact desired orientation may be challenging, making trade-offs necessary.
-However, when feasible, it is preferable to use more links to better satisfy the orientation requirement – thus, $f_{2}$ takes precedence over $f_{3}$, and $f_{1}$ is given overall priority.
-Based on this, the optimization problem becomes to reach the target’s pose using the least amount of resources.
-Therefore, we should first verify that the configuration being examined can reach the target ($f_{1}$), and then proceed to minimize its components.
-It is important to note that although the robot must precisely reach the target, $f_{1}$ should not be treated as a constraint (i.e., solutions with a non-zero tip-to-target distance should not be deemed infeasible).
-If $f_{1}$ is treated as a constraint, the search algorithm may lose its heuristic effectiveness, leading to random convergence [^4] This claim is debatable, and it depends on the constraint-handling method employed: it would hold true for death or static penalty methods, but it could be valuable to explore results using dynamic methods or separation-based approaches [84], despite the higher risk of obtaining infeasible solutions only..
-Using methods that retrieve Pareto optimal solutions, known as generating methods [85], may not be ideal for our situation unless we intentionally limit part of the objective space.
-Instead, we aim to integrate the three objectives into a single function, referred to as a fitness function.
-In addition to prioritizing the first objective, single-objective optimization methods generally have lower computational complexities than multi-objective methods.
-To solve this design optimization problem, we employ the Rank Partitioning method, which was introduced in the optimizer for the 2D case [67].
-Rank partitioning is a non-parametric survival strategy for Multi-Objective EC algorithms (or generically for population-based optimization methods) that ranks each individual based on a predefined order of the objectives.
-It sorts the population for the first objective (i.e., the one with the highest priority) and then creates partitions based on their objective rank; then, it sorts each partition for the second objective, and so on.
-Finally, it returns only one solution: the optimal based on the given order of priorities.
-It can be applied to multi or many-objective optimization (i.e., with more than three objectives [86]).
-3.3.1 Solving the Inverse Kinematics with the Desired Orientation
-The first two objectives are to ensure that the end effector of each configuration reaches the respective target ($f_{1}$) with the desired orientation ($f_{2}$).
-This is a well-known problem of inverse kinematics [87], which can be solved by minimizing the distance between the end effector and the target for each configuration.
-This is usually solved in task space (i.e., Cartesian coordinates in the environment) rather than in joint space (i.e., polar coordinates defining the values of each robotic joint).
-Therefore, since the configurations $\zeta$ in Eqn.
-(1) are defined in polar coordinates ($\zeta^{p}$), it is necessary to transform it into its corresponding configuration in Cartesian coordinates ($\zeta^{c}$) through forward kinematics as in Eqn.
+3.2 Fitness Evaluation and Dynamic Genotype Change
+We included the following objective functions:
+  • 1.
+    $f_{1}$: For each configuration, minimize the distance between the end effector and the target – inverse kinematic solver.
+  • 2.
+    $f_{2}$: For each configuration, minimize the displacement between the desired orientation segment and the robot’s last link – inverse kinematic solver.
+    This is done by transforming configurations $\zeta$ in Eqn.
+(1) from polar coordinates ($\zeta^{p}$, i.e., in joint space) to Cartesian coordinates ($\zeta^{c}$, i.e., in task space) through forward kinematics as in Eqn.
 (4) – details are reported in Section 1.1 of the Supplementary Material.
-The configuration $\zeta^{c}$ is composed of $n+1$ points – i.e., all nodes plus the end effector.
+The configuration $\zeta^{c}$ includes the end effector, resulting in $n+1$ nodes.
     $\zeta^{p}=\begin{bmatrix}\theta^{(x)}_{1}&\theta^{(y)}_{1}&l_{1}\\ \vdots&\vdots&\vdots\\ \theta^{(x)}_{n}&\theta^{(y)}_{n}&l_{n}\\ \end{bmatrix}\quad\xrightarrow[kinematics]{forward}\quad\zeta^{c}=\begin{bmatrix}x_{1}&y_{1}&z_{1}\\ \vdots&\vdots&\vdots\\ x_{n+1}&y_{n+1}&z_{n+1}\end{bmatrix}$ (4)
-Figure 6: The phenotype of a robot generated with the genotype described in Eqn.
-(2), having $n=10$ links.
-In this case, only five are actually active to reach the target correctly ($\bar{n}=5$).
-Moreover, the last link is not fully extended, unlike the internal links that make up the rest of the body: this information is codified by extending the genotype as in Eqn.
-(6).
-The five links after the end effector (represented by dotted lines) do not need to be included in the design $\delta$.
+    • (a)
+      $f_{1-2}$: We merged these two objectives into a single one, which is to minimize the distance between the closest node to the target’s orientation segment and the segment itself, allowing the robot to correctly aim for the target’s pose as shown in Fig.
+5(a).
+This is calculated according to Eqn.
+(5) – the distance from a node to the $i$-th target’s orientation segment $s_{i}$ is reported in Algorithm S.3 of the Supplementary Material.
+      $f_{1-2}=\sum_{i=1}^{|t|}{\min_{\forall j\in n}{(\texttt{pointSegmentDistance}(\zeta^{c}_{i,j+1},s_{i})})}$ (5)
+      This formulation requires some additional information to the basic genotype described in Eqn.
+(2): once the closest node to the target’s orientation segment is found (at index $\epsilon$), the remainder of the configuration’s link is unused (the dotted links in Fig.
+5(a)), and we replace them with links that go straight to the target.
+This results in a different orientation from the angles specified for the $\epsilon$-th joint (defined as $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$), a number of actually used links in the overall configuration ($\bar{n}$), and the length of the last link that does not necessarily need to be fully unfolded ($l_{\bar{n}}$).
+The procedures for retrieving these values are reported in the Supplementary Material.
     (a) Inverse Kinematic Solver
     (b) Phenotype Description
-Figure 7: (a) When the inverse kinematic solver minimizes the distance node-target, there is no guarantee for the robot to have the last link aligned with the desired target orientation; however, when the solver minimizes the distance node-orientation, the robot will reach the desired orientation and can simply grow straight to the target after that.
+    Figure 5: (a) When the inverse kinematic solver minimizes the distance node-target $d$, there is no guarantee for the robot to have the last link aligned with the desired target orientation; however, when the solver minimizes the distance node-orientation, the robot will reach the desired orientation and can simply grow straight to the target after that.
 (b) The phenotype codified by a solution $\psi$ ignores every link after $\epsilon$ originally generated in the genotype and aims directly to the target for the remaining $\bar{n}-\epsilon$ links.
 The last link might not be fully everted, with a length smaller than the one originally generated in the genotype – indicated with the term $l_{\bar{n}}$.
-The first issue to be addressed is the location of the end effector in a configuration: the population of any EC algorithm will be filled with random values for each solution $\psi$, meaning that the robotic configurations will explore the environment randomly to eventually converge with the evolutionary metaheuristic.
-Fig.
-6 shows that the tip of the robot, or the end effector, can be placed anywhere on any link as the robot grows along its configuration.
-The link whose endpoint corresponds to the end effector is indicated with the index of $\bar{n}$.
-In case $n\neq\bar{n}$, in which $\bar{n}$ links are active and the rest $n-\bar{n}$ are folded within the robot, it is possible that the $\bar{n}$-th link is only partially everted – i.e., the target is reached before the link is fully grown.
-Based on the definition in Eqn.
-(2), multiple configurations share the same design because the robot can take different configurations with different angles, but the link lengths cannot change.
-However, when different targets require a different amount of growth, as shown in Fig.
-4(b), it is necessary to modify the value of the last link for each configuration (i.e., $\bar{n}$) such that the link length only takes into account the active part of the link.
-In that case, the $\bar{n}$-th link length is not taken from the design $\delta$ (i.e., the last row of $\psi$ in Eqn.
-(2)), but is defined as the distance between the $\bar{n}$-th node and the target, and indicated with the term $l_{\bar{n}}$.
-The second issue to be addressed is to guarantee that the configuration’s last link(s) lie(s) on the target’s desired orientation segment.
-The left-hand side of Fig.
-7(a) shows that when minimizing the distance node $\bar{n}$ to target, the optimized configuration might end up not being aligned with the desired orientation.
-Rather than also minimizing the difference between the orientation of the robot’s last link and the desired one, it is more efficient to combine both objectives $f_{1}$ and $f_{2}$ into a single one, namely $f_{1-2}$.
-The right-hand side of Fig.
-7(a) shows an alternative minimizing metric: reducing the distance between the end effector $\bar{n}+1$ to the target’s desired orientation segment.
-With this formulation, once the distance is minimized, the robotic configuration will intersect with the target’s orientation segment at the end effector $\bar{n}+1$.
-However, the end effector would still not be on the target, but somewhere on the segment: at this point, the robot can simply evert toward the target to eventually reach it – as shown in Fig.
-7(b).
-This requires additional information to be stored in the genotype of a solution $\psi$.
-Firstly, the node selected to reach the target’s orientation segment is not the end effector but the closest node to the segment (at index $\epsilon$) among all the configuration nodes.
-Secondly, the rotation angles at the $\epsilon$-th joint (namely, $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$) must not follow the ones indicated in the configuration $\zeta$, but should be geometrically calculated such that the robot is oriented in the direction of the target (the procedure is reported in Sections 1.2 and 1.3 of the supplementary material).
-Fig.
-7(b) shows the phenotype described by the genotype with the new extra information.
-Considering these two issues, the additional information needs to be codified in the data structure/genotype of solutions $\psi$, whose elements define the decision variables of the optimization problem.
-The matrix $\psi$ then is defined as in Eqn.
-(6), where the left submatrix is the solution genotype $\psi$ defined in Eqn.
-(2), and the right submatrix is composed of the extra information for each configuration.
-Specifically, for each configuration, every pair of rows of the right submatrix contains:
-  • 1.
-    $\epsilon$, the index of the closest node to the target’s orientation segment;
-  • 2.
-    $\theta^{(x)}_{\epsilon}$, the angle on the x-axis for the $\epsilon$-th link of the robot to be aligned to the target’s orientation segment;
   • 3.
-    $\theta^{(y)}_{\epsilon}$, the angle on the y-axis for the $\epsilon$-th link of the robot to be aligned to the target’s orientation segment;
-  • 4.
-    $\bar{n}$, the number of active nodes/links for each configuration (with $\bar{n}\leq n$) – namely the index of the last active node of a configuration; and
-  • 5.
-    $l_{\bar{n}}$, the length of the last link for each configuration, right before the end effector.
-Note that, to keep a matrix shape, some elements are set to zero as they are unused: (i) the second row of each configuration, which defines rotations on the y-axis, only contains the last angle rotation $\theta^{(y)}_{\epsilon}$; and (ii) the last row, which defines the link lengths, are all fixed for a single design and do not need any extra information.
-[…]
-While the left submatrix is either randomly generated in the first generation of the EC algorithm or generated by variation when producing offspring, the right submatrix is filled during solution evaluation based on the geometrical requirements needed to reach each specific target.
-Also, it is important to note that the link lengths obtained in the design $\delta$ in Eqn.
-(3) serve as a reference only for the target farthest from the home base $h$, and that different configurations may use partial information from $\delta$ based on the number of active links.
-The combined fitness $f_{1-2}$ is expressed as in Eqn.
-(7), where the distance between a node of the robot and the segment $s_{i}$ is calculated with the point-segment distance (introduced by Stroppa et al.
-[88] for 2D and extended for this work’s purposes in 3D as reported in Algorithm S.3 of the supplementary material).
-This computes the distance from the point to its projection onto the segment or to the nearest edge.
-This formulation also retrieves the index $\epsilon$ of the node that will intersect with the target’s orientation segment.
-$f_{1-2}=\sum_{i=1}^{|t|}{\min_{\forall j\in n}{(\texttt{pointSegmentDistance}(\zeta^{c}_{i,j+1},s_{i})})}$ (7)
-3.3.2 Optimizing the Kinematic Components of the Robot
-This part is the same as the 2D case presented in the previous work [67], except for having two rotational angles for each node; however, we report it fully for a comprehensive description and reproducibility.
-The objective functions $f_{1}$ and $f_{2}$ (inverse kinematics) take precedence over $f_{3}$ (optimal components).
-After combining the first two into a single fitness function $f_{1-2}$, the next step is penalizing excessively long solutions with too many nodes.
-Ideally, we aim for a robot that gradually aligns with the target’s orientation segment, creating a smooth curve.
-The robot depicted in the right-hand side of Fig.
-5(b) is a better solution compared to the one shown in Fig.
-5(a): it not only positions all nodes closer to the orientation segment but also uses fewer nodes to reach the target.
-Objective $f_{3}$ focuses on optimizing the robot’s components to minimize manufacturing time and resources and cannot be captured by a single property.
-Thus, we treat it as a macro objective composed of the following:
-  • 1.
-    $f_{3.1}$: minimize the number of links of each configuration to reduce redundant DoFs, as in Eqn. (8);
-    $f_{3.1}=\sum_{i=1}^{|t|}\bar{n}_{i}$ (8)
-  • 2.
-    $f_{3.2}$: minimize the overall length of the robot to promote shorter configurations, as in Eqn. (9); and
+    $f_{3}$: For each configuration, minimize the robot’s components to promote a quicker and more cost-effective manufacture. We split this objective into the following:
+    • (a)
+      $f_{3.1a}$: For each configuration, minimize the number of links needed to reach the target’s orientation segment, as in Eqn. (6).
+      $f_{3.1a}=\sum_{i=1}^{|t|}(\epsilon_{i}-1)$ (6)
+    • (b)
+      $f_{3.1b}$: For each configuration, minimize the number of links lying on the target’s orientation segment, as in Eqn. (7).
+      $f_{3.1b}=\sum_{i=1}^{|t|}(\bar{n}_{i}-\left(\epsilon_{i}-1\right))$ (7)
+    • (c)
+      $f_{3.2}$: Minimize the overall length of the robot, which is the longest robot’s length among all configurations, as in Eqn. (8).
       $f_{3.2}={\max_{\forall i\in|t|}{\left(\left(\sum_{j=1}^{\bar{n}_{i}-1}{l_{j}}\right)+l_{\bar{n}_{i}}\right)}}$ (8)
-  • 3.
-    $f_{3.3}$: the change in direction between links in all configurations, or undulation, to favor trajectories that are non-intersecting and not “wavy”, calculated as a percentage of the total length of all configurations, as shown in Eqn.
+      (a) Wavy robot
+      (b) Trade-off nodes/orientation
+      Figure 6: (a) A poorly designed robot featuring a wavy configuration.
+(b) Reducing the number of links and increasing their length helps solve this issue, a trade-off is necessary to make sure the robot smoothly reaches the target.
+    • (d)
+      $f_{3.3}$: For each configuration, minimize the directional change of the links, or undulation, to avoid “wavy” configurations are shown in Fig.
+6.
+This is evaluated in percentage over the number of links, as shown in Eqn.
 (9) – note that the formula checks the change in the angle between the $j$-th and $(j+1)$-th link in both the $xz$ and $yz$-plane projections ($\Delta\theta^{(xz)}$ and $\Delta\theta^{(yz)}$, respectively).
       $f_{3.3}=\frac{\sum_{i=1}^{|t|}{\left(\frac{\forall k\in[xz,yz],\;\sum_{j=1}^{\epsilon_{i}-1}{\left[\operatorname{sign}{(\theta^{(k)}_{i,j})}\neq\operatorname{sign}{(\theta^{(k)}_{i,j+1})}\right]}}{2\cdot\epsilon_{i}}\right)}}{|t|}\cdot 100$ (9)
-The priorities among these three sub-objectives can be adjusted to refine our preference-based method, potentially eliminating the need for weight parameters.
-Although it might appear that the number of links and the total length of the robot ($f_{3.1}$ and $f_{3.2}$, respectively) have a linear relationship (i.e., more links lead to a longer robot), this is not necessarily true.
-A robot with many small links can be shorter than one with fewer, very long links.
-We prefer robots with fewer links because the time required to manufacture a link is not affected by its length ($f_{3.1}$ takes precedence over $f_{3.2}$).
-On the other hand, the number of links and their orientations ($f_{3.1}$ and $f_{3.3}$, respectively) affect the robot’s shape differently.
-A robot with many links is more likely to be wavy, as illustrated in Fig.
-5(a).
-Whether high-DoF non-wavy robots are preferable to low-DoF wavy ones may not be straightforward and could depend on the specific task (e.g., navigating a maze might require varying directions); however, we found that fewer links are preferable for manufacturing ($f_{3.1}$ takes precedence over $f_{3.3}$), but other manufacturers might have different priorities and choose to reverse this order.
-Finally, the undulation of a robot and its length ($f_{3.3}$ and $f_{3.2}$, respectively) are strongly correlated.
-A robot that directly targets the destination will generally be shorter than one that moves through the environment before reaching the target, as demonstrated in Fig.
-5(a).
-Therefore, we consider it more important to construct non-wavy robots, which will ultimately lead to shorter configurations ($f_{3.3}$ takes precedence over $f_{3.2}$).
-Furthermore, we found that to enhance the quality of the solutions obtained, the objective $f_{3.1}$ should be broken down into two sub-objectives:
+Since this evaluation dynamically adds extra information to the genotype, it is necessary to extend the solution matrix $\psi$ by including, for each configuration:
   • 1.
-    $f_{3.1a}$: minimize the number of links to reach the target’s orientation segment, as in Eqn. (11); and
-    $f_{3.1a}=\sum_{i=1}^{|t|}(\epsilon_{i}-1)$ (11)
+    $\epsilon$, the index of the configuration node/joint that is closest to the target’s orientation segment;
   • 2.
-    $f_{3.1b}$: positioned along the target’s orientation segment, as in Eqn. (12).
-    $f_{3.1b}=\sum_{i=1}^{|t|}(\bar{n}_{i}-\left(\epsilon_{i}-1\right))$ (12)
-By making this division, we specifically direct how the configuration should be arranged rather than just minimizing the total number of links across different configurations.
-This involves prioritizing the reduction of links required to align with the target’s orientation segment.
-Once this segment is reached, we then focus on minimizing the number of links needed to reach the target itself.
-Note that the final step can be effectively achieved by manually maximizing the lengths of these links.
-Although the optimization algorithm should ideally learn this heuristic, practical implementation might require some manual adjustments due to the extensive search space.
-3.4 Separation of the Objectives with Rank Partitioning
-1
-input : $N$, population size.
-input : $F$, a matrix $N\times 9$ having the following information for each individual: fitness for each objective (integer values first followed by real values), fitness after partitioning ($rank$), reference to the individual in the population ($ref$).
-A single row of $F$ is: [(1)[(1)[ ( 1 ): $F_{1-2\mathbb{N}}$, $(2)$: $F_{3.1a}$, $(3)$: $F_{3.3}$, $(4)$: $F_{3.1b}$, $(5)$: $F_{3.2\mathbb{N}}$, $(6)$: $F_{1-2\mathbb{R}}$, $(7)$: $F_{3.2\mathbb{R}}$, $(8)$: $rank$, $(9)$: $ref]_{i}$ $\forall i\in[1,N]$.
-input : $binSize$, distance between partitions for each real-coded objective (i.e., one for $F_{1-2}$ and one for $F_{3.2}$).
-output : The updated matrix $F$ with the value of $rank$ after partitioning.
-2
-3begin
-4       for $i\leftarrow 1$ to $N$ do
-5             $F_{i,1}\leftarrow F_{i,6}-\textnormal{{mod}}(F_{i,6},binSize_{1-2})$
-6             $F_{i,5}\leftarrow F_{i,7}-\textnormal{{mod}}(F_{i,7},binSize_{3.2})$
-7
-8
-9      sortrows($f,>$) on column 1-5 for all rows
-10       $d\leftarrow 0_{N\times 5}$
-11       for $i\leftarrow 2$ to $N$ do
-12             for $j\leftarrow 1$ to $5$ do
-13                   $d_{i,j}\leftarrow\textnormal{{abs}}(F_{i,j}-F_{i-1,j})$
-14
-15
-16
-17      $start\leftarrow 1$
-18       for $i\leftarrow 1$ to $N$ do
-19             for $j\leftarrow 1$ to $5$ do
-20                   $partLimit\leftarrow partLimit+d_{i,j}$
-21
-22            if $partLimit>0$ then
-23                   $end\leftarrow i$
-24                   sortrows($f,>$) on column 6-7 from row start to end-1
-25                   $start\leftarrow end$
-26
-27
-28
-29      for $i\leftarrow 1$ to $N$ do
-30             $F_{i,8}\leftarrow i$
-31
-32      return $F$
-33
-Algorithm 1 Rank Partitioning [67]
-Based on the observations about objective priorities, we employ a strategy that prioritizes solutions with fewer, well-structured components, but only if they are sufficiently close to the target’s orientation segment.
-The Rank Partitioning algorithm was introduced in the previous work by Stroppa [67] and is used to retrieve the best solution of a MOOP given a predetermined order of priorities among objectives, without the burden of compromising the optimality of the highest prioritized objectives and choosing weights for each of them (i.e., unlike weighted sum methods [78, 89]).
-Rank partitioning divides the population into $k$ partitions (bins) with fixed intervals (bin sizes) according to their kinematic fitness $f_{1-2}$.
-Individuals with better fitness are assigned to the first segments, while those with worse fitness are placed in the last segments – in this specific context, where all objectives will be minimized, the best fitness is the one with the lowest value.
-This is repeated for each objective: the partitions made for the first objective will be divided into more partitions based on the second objective, and so on, until all the objectives are partitioned.
-The key idea is that individuals within the same partition are considered equally optimal in terms of the partitioned objective so that the best among them can be selected based on the following objectives.
-Just like in the 2D scenario [67], we applied Rank Partitioning in the following order for each partition:
-  1. 1.
-    distance from the target/inverse kinematics ($f_{1-2}$);
-  2. 2.
-    number of links to reach the targets’ orientation segment ($f_{3.1a}$);
-  3. 3.
-    percentage of undulation among configurations ($f_{3.3}$);
-  4. 4.
-    number of links on the targets’ orientation segment ($f_{3.1b}$); and
-  5. 5.
-    length of the robot/final design ($f_{3.2}$).
-This approach works only when ties happen within partitions, leading to nested sub-partitions that follow the prioritized order of objectives.
-This creates a challenge when handling objectives with continuous domains, as we deal with ranges or bins of fixed intervals rather than individual values (similar to a histogram).
-To address the issue, Rank Partitioning duplicates all continuous objectives (i.e., $f_{1-2}$ and $f_{3.2}$) by adding a further objective composed by the lower-bound discrete value of their bin.
-Discrete objectives can be used with no modification (i.e., $f_{3.1}\in[1,n\cdot|t|]$ and $f_{3.3}\in[0,100]$).
-In this manner, the partitioning operation can be resolved using a straightforward sorting algorithm that accounts for ties.
-When a tie occurs in the objective with the lowest priority, individuals are ranked based on their original continuous values of continuous-domain objectives (i.e., $f_{1-2}$ and $f_{3.2}$), which should be preserved rather than substituted by their discrete bin values.
-Finally, Rank Partitioning updates the fitness value by assigning a rank to each individual in the population, which increases linearly according to their position within the partitions (i.e., sequentially numbering each individual after the partitioning process from 1 to $N$, with $N$ being the population size).
-Algorithm 1 shows Rank Partitioning on the soft growing robot problem, which runs in $\mathcal{O}(N\texttt{log}N)$, where $N$ represents the population size.
-The input $F$ is a “fitness” matrix $N\times 9$ containing, for each individual, (i) its kinematic fitness $f_{1-2}$, (ii) the four sub-objectives describing the components $f_{3}$, (iii) the original values of the continuous objectives that were evaluated by their lower-bound bin value, (iv) the updated rank value, calculated by the procedure, will be utilized by the selection operators for the reproduction and survival of individuals, and (v) a reference to the individual in the population [^5] The reference is used to link fitness in the matrix and the corresponding individual in the population since only the fitness matrix will be sorted to save computational time.
-This is simply a pointer to the respective individual in the population..
-Lines 2-4 convert continuous objectives (denoted by the subscript $\mathbb{R}$) into discrete values (denoted by subscripts $\mathbb{N}$) based on the specified bin size [^6] Note that the subscript $\mathbb{N}$ is just a placeholder since these values are not limited to natural numbers only as their bin size can be smaller than 1..
-Line 5 sorts the rows of the fitness matrix according to the five discrete objectives in the order provided.
-Lines 6-9 determine partition boundaries by computing differences between consecutive elements in each sorted column of the fitness matrix: if two consecutive elements in the same column are equal, they belong to the same partition.
-Summing these differences for each row (lines 12-13) helps the algorithm identify partition boundaries.
-Lines 11-17 re-sort each row by the original continuous values, but only within partitions, ensuring the order from line 5 remains intact.
-Finally, lines 18-20 assign the updated rank to each individual.
-In multi-objective optimization, this approach acts as a restriction within the objective space, focusing only on the segment of the Pareto optimal front with the lowest values for one objective: the first partition represents the optimal solution for $f_{1-2}$, allowing us to select the individual that minimizes $f_{3}$ as the best fit.
-The key benefits of this method are: (i) it enforces priority on a specific objective (e.g., inverse kinematics in our case), (ii) it eliminates the need to select weight values as in typical preference-based methods, and (iii) it provides flexibility for future research to incorporate additional sub-objectives that may be relevant but were not considered in this study (i.e., for tackling many-objective optimization problems).
+    $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$, the angles on the x and y-axes defining the orientation for the $\epsilon$-th link, dynamically set to go in the direction of the target;
+  • 3.
+    $\bar{n}$, the number of active nodes/links (with $\bar{n}\leq n$) – namely the index of the last active node of a configuration; and
+  • 4.
+    $l_{\bar{n}}$, the last link’s length.
+The complete matrix $\psi$ is defined as in Eqn.
+(10), and the respective phenotype that it represents is shown in Fig.
+5(b) – example reports one configuration only for the sake of clarity.
+Note that, being a matrix, some elements are set to zero as they are unused: (i) the second row of each configuration, which defines rotations on the y-axis, only contains the last angle rotation $\theta^{(y)}_{\epsilon}$; and (ii) the last row, which defines the lengths of the links, is fixed for a design and do not need any extra information.
+[…]
+Figure 7: The genotype described in Eqn.
+(2) yields the phenotype in the example, with a budget of $n=10$ links.
+However, only five of them are required to correctly reach the target ($\bar{n}=5$), an information included in the genotype of Eqn.
+(10).
 3.3 Constraints
 The problem’s constraints are similar to the 2D case presented in the previous work [67], with the respective additions of one more dimension for the rotational angles and one more constraint on obstacle collision that was not previously – and erroneously – considered.
-Every configuration in the solutions is subject to eight kinematic constraints in addition to the bounds on decision variables (i.e., $\Delta_{\theta}$ and $\Delta_{l}$, which are parameters the user sets prior to manufacturing).
-Since the final links (i.e., from the $\epsilon$-th to the $\bar{n}$-th node) are automatically determined to reach the target as described in Section 3.3.1, the orientation angles $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$ between the last active link and the target’s orientation segment depends on the kinematic chain produced by the genetic operators.
-As a result, it may exceed the bounds $\Delta_{\theta}$, as illustrated in Fig.
-8(a).
-We defined four inequality constraints, as shown in Eqn.
-(13), to mark solutions as infeasible when configurations have $\theta^{(x)}_{\epsilon}$ and $\theta^{(y)}_{\epsilon}$ beyond these limits.
+Every configuration in the solutions is subject to eight kinematic constraints (depicted in Fig.
+8) in addition to the decision-variable bounds described in Section 3.1.
+Figure 8: Examples of solutions violating constraints – i.e., infeasible.
+(a) Excessive steering when reaching the target; (b) a small final link that makes it impossible for the robot to carry a gripping mechanism; (c) overshooting the target, resulting in excessive angles; (d) an obstacle collision; and (e) a collision non detected in the configuration but as result of motion, which requires a link to be everted before steering.
+The first four constraints concern the angles at the $\epsilon$-th node, which are set by the evaluation function such that, from that point on, the robot grows in the direction of the target – as described in Section 3.2.
+Since these angles are geometrically calculated, there is no guarantee that they respect the bounds $\Delta_{\theta}$, as illustrated in Fig.
+8(a); therefore, we included the four inequality constraints defined in Eqn.
+(11).
 Note that the bound $\Delta_{\theta}$ can also differ between axes, but for the sake of simplicity, we will report it as the same value.
-Figure 8: Some examples of infeasible solutions.
-In (a), excessive steering when reaching the target; in (b), a very small link at the $\epsilon$-th link right after steering that cannot accommodate a gripper device; in (c), not reaching the target with the desired orientation; in (d), an obstacle collision; and in (e) a collision non detected in the configuration but as result of the motion planner, which requires a link to be everted before steering.
 $\theta^{(x)}_{\epsilon}\geq{\Delta_{\theta}}^{(L)}\wedge\theta^{(x)}_{\epsilon}\leq{\Delta_{\theta}}^{(U)}$ (11)
 $\theta^{(y)}_{\epsilon}\geq{\Delta_{\theta}}^{(L)}\wedge\theta^{(y)}_{\epsilon}\leq{\Delta_{\theta}}^{(U)}$
-Similarly, the length of each configuration is determined by the distance between the $i$-th target and the $\bar{n}_{i}$-th node, which may result in configurations where the length of the final link, $l_{\bar{n}}$, falls below its lower bound limit.
-In practice, this becomes an issue only when the final link needs to adjust its direction to align with the target’s orientation segment.
-Considering that these manipulators are likely equipped with a gripping mechanism at the tip (e.g., the tip mount proposed by Jeong et al.
+The fifth constraint considers the case in which the length of the last link is smaller than the lower bound ${\Delta_{l}}^{(L)}$, as shown in Fig.
+8(b).
+This is an issue only when the final link needs to adjust its direction to align with the target’s orientation segment – i.e., there is only one link in on the segment, or $f_{3.1b}=1$, Eqn.
+(7): these soft growing robots are likely equipped with a gripper device that moves with the end effector along the robot’s body (e.g., the tip mount proposed by Jeong et al.
 [81] or the mechanism to prevent buckling proposed by Coad et al.
-[82]), it is essential to account for the necessary space to position the gripper properly on the target’s orientation segment for accurate grasping.
-We formulated the bounds ${\Delta_{l}}^{(L)}$ to account for this constraint.
-As a result, any solutions featuring configurations where a single link on the target’s orientation segment ($f_{3.1b}=1$, Eqn.
-(12)) is insufficiently long are considered infeasible, as illustrated in Fig.
-8(b).
-This leads to the definition of a fifth inequality constraint, as expressed in Eqn.
+[82]).
+For this reason, it is essential to account for the necessary space to position this gripper properly on the target’s orientation segment for accurate grasping.
+The respective inequality constraint is expressed in Eqn.
 (12).
 $\text{if}\;f_{3.1b}=1\text{,}\quad l_{\bar{n}}\geq{\Delta_{l}}^{(L)}$ (12)
-The $\epsilon$-th node, identified as the nearest joint to the target’s orientation segment, is determined using the point-segment distance method (details are reported in Section 1.2 of the supplementary material).
+The sixth constraint handles an edge case of the point-segment distance method of Eqn.
+(5) – details of the procedure are reported in Section 1.2 of the Supplementary Material.
 Since this process may designate one of the segment’s endpoint as the nearest point to $\epsilon$ – rather than $\epsilon$’s projection on the segment – if that endpoint corresponds to the target, the resulting solution would resemble the configuration shown in Fig.
@@
 Lastly, we impose two constraints related to obstacle collisions.
-If the environment contains elements that are not targets, they must be treated as obstacles.
-Any solution that results in a collision with these obstacles is marked as infeasible, as illustrated in Fig.
+The first one imposes that no collision is generated in the solution, as a robot colliding with an obstacle would be an infeasible solution – as illustrated in Fig.
 8(d).
-In our optimization problem, obstacle collisions are treated as an equality constraint, formulated as shown in Eqn.
-(16).
-$\texttt{intersections}(\psi,o)=0$ (16)
-The function in Eqn.
-(16) calculates the number of intersections between each obstacle in the set $o$ and every link of each configuration in $\psi$.
+The equality constraint in Eqn.
+(14) evaluates the intersections between the links in $\psi$ and the obstacles in the set $o$.
 The full procedure is reported in Section 1.4 of the Supplementary Material, and it operates with a complexity of $\mathcal{O}(\overline{n}\cdot|o|\cdot|t|)$.
@@
 More complex obstacles can be modeled by combining multiple cylinders: a pillar is represented by a single cylinder; a wall by a line of cylinders; and a hole in a wall by stacking cylinders in a bottom layer, a top layer, and two taller cylinders at the sides, forming a hole in the middle.
-During problem setup, users can select a radius slightly larger than that of the obstacle’s circumscribed cylinder to prevent contact between the robot and the obstacle.
+During problem setup, users can select a larger radius than that of the obstacle’s circumscribed cylinder to prevent contact between the obstacle and the manipulator.
 Note that avoiding obstacles might also be considered as another objective function: one can (i) minimize obstacle collisions or (ii) maximize the distance between any robot link and obstacles.
@@
 Therefore, considering obstacle collisions as a constraint is the best option for our optimization problem.
+$\texttt{intersections}(\psi,o)=0$ (14)
 The last constraint regards a possible obstacle collision when actuating the manipulator – i.e., the static configuration $\zeta_{i}$ does not collide with obstacles, but the actuation to move to $\zeta_{i}$ from any other configuration $\zeta_{j}$ will cause a collision.
@@
 8(e) shows an example: if the configuration requires the robot to grow of a length equal to $\Delta^{(L)}_{l}$ within the obstacle, then we have a collision – note that the figure represents a projection in 2D of a spherical cone.
-Note that a motion planner could actually find a feasible alternative path without colliding with the obstacle (i.e., by bending one of the previous joints); however, checking this during optimization would require running the motion planner [83], which would require too much in terms of computational resources.
+A motion planner could actually find a feasible alternative path without colliding with the obstacle (i.e., by bending one of the previous joints); however, checking this during optimization would require running the motion planner [83], which would require too much in terms of computational resources.
 Therefore, we simply mark a configuration like this one as infeasible.
@@
 $\texttt{cannotSteer}(\psi,o)=0$ (15)
-4 Experimental Results and Discussion
-This section illustrates the experiments to (i) assess whether the optimizer can retrieve optimal solutions on tasks with large dimensionality and many obstacles; (ii) determine which EC algorithm is the best fit for solving our design problem among four different options; and (iii) evaluate the robustness of the Rank Partitioning algorithm when included in the loop of different EC algorithms.
+4 Experiments
+We ran a set of experiments to (i) assess whether the optimizer can retrieve optimal solutions on tasks with large dimensionality and many obstacles; (ii) determine which EC algorithm is the best fit for solving our design problem among four different options; and (iii) evaluate the robustness of the Rank Partitioning algorithm when included in the loop of different EC algorithms.
+4.1 Optimization Methods
+To solve this design optimization problem, we employ the Rank Partitioning method, which was introduced in the optimizer for the 2D case [67].
+Rank partitioning is a non-parametric survival strategy for multi-objective EC algorithms (or generically for population-based optimization methods) that ranks each individual based on a predefined order of the objectives.
+It sorts the population for the first objective (i.e., the one with the highest priority) and then creates partitions based on their objective rank; then, it sorts each partition for the second objective, and so on.
+Finally, it returns only one solution: the optimal based on the given order of priorities – specifically, the priority order we selected for our objectives is $f_{1-2}$ (to ensure the robot reaches the targets), $f_{3.1a}$ (to avoid configurations roaming before reaching the correct orientation), $f_{3.2}$ (to prevent wavy configurations), $f_{3.1b}$ and $f_{3.2}$ (to minimize material waste).
+The advantages of such a method are: (i) it allows us to retrieve the optimal solution based on our preference without compromising trade-offs; (ii) there is no need to retrieve the full Pareto optimal front, saving computational resources; and (iii) it outperforms weighted sum based methods while without featuring their disadvantages [78, 89].
+In this section, we seek evidence for a fourth advantage: this method can be applied to any EC algorithm, or any multi or many-objective optimization (i.e., with more than three objectives [86]).
+More details about Rank Partitioning are reported in Section 2 of the Supplementary Material or in our previous work [67].
 input : The optimization problem $op$ (in our case, described by $t$, $h$, $o$, $n$, $\Delta_{\theta}$, $\Delta_{l}$) including a set of constraints $c$
@@
 Algorithm 1 EC framework with Rank Partitioning
-4.1 Optimization Methods
 We tested the design optimizer on four different EC algorithms: Genetic Algorithm (GA) [91], Particle Swarm Optimization (PSO) [92], Differential Evolution (DE) [93], and Big Bang-Big Crunch algorithm (BBBC) [94].
-Each of these EC algorithm has its own parameters, which are described in the following subsections; however, some of the parameters are shared as they all comply with the basic framework of EC algorithms, reported in Algorithm 2: an initial population $pop$ is initialized with $N$ random individuals (line 2), evaluated returning the fitness of each individual $F^{pop}$ penalized in case of constraint violation (line 3), ranked based on the Rank Partitioning algorithm (line 4, runs Algorithm 1), and then it undergoes the generational loop for $G$ iterations (line 5-14) in which variation operators generate a new population of offspring $off$ (line 6), evaluate them returning their fitness $F^{off}$ with penalty in case of constraint violation (line 7), and select which individuals among $pop$ and $off$ will survive and be part of the next generation based on the survival scheme (lines 8-14) – specifically, when elitism is applied, the Rank Partitioning algorithm is executed on both populations (line 9), otherwise only on the offspring (line 12).
+Each of these EC algorithm has its own parameters, which are described in the following subsections; however, some of the parameters are shared as they all comply with the basic framework of EC algorithms, reported in Algorithm 1: an initial population $pop$ is initialized with $N$ random individuals (line 2), evaluated returning the fitness of each individual $F^{pop}$ penalized in case of constraint violation (line 3), ranked based on the Rank Partitioning algorithm (line 4, runs Algorithm S.11 of the Supplementary Material), and then it undergoes the generational loop for $G$ iterations (line 5-14) in which variation operators generate a new population of offspring $off$ (line 6), evaluate them returning their fitness $F^{off}$ with penalty in case of constraint violation (line 7), and select which individuals among $pop$ and $off$ will survive and be part of the next generation based on the survival scheme (lines 8-14) – specifically, when elitism is applied, the Rank Partitioning algorithm is executed on both populations (line 9), otherwise only on the offspring (line 12).
 Note that, with Rank Partitioning, the best individual to be saved and returned by the method is not the individual with minimum/maximum fitness value, but it is the one having rank $=1$ after executing Rank Partitioning between the previous best and the best of the current generation (line 15) – regardless on whether the method is elitist or not.
@@
 This process enables particles to converge toward optimal or near-optimal solutions while avoiding getting trapped in local optima, making PSO particularly effective for solving complex, nonlinear optimization problems.
-Table 3: Experimental conditions and parameters (PSO)
+Table 4: Experimental conditions and parameters (shared among EC algorithms)
 Parameter Value
-Inertia Weight ($\omega$) $0.6,0.8,1.0$ [101]
-Cognitive Constant ($c_{1}$) $0.5,2.5,5.0$ [102, 103]
-Social Constant ($c_{2}$) $0.5,2.5,5.0$ [102, 103]
-Table 3 reports the parameter settings we used for PSO.
+Population Size ($N$) $500$
+Max Number of Generations ($G$) $500$
+Constraint Handling Static Penalty [95, 84] $R=100$
+Rank Partitioning Bin Size for $f_{1-2}=0.5$ for $f_{3.2}=5.0$
+Table LABEL:tab:param_pso reports the parameter settings we used for PSO.
 The inertia weight $\omega$ is between [0.5, 1.0] after the random $\omega$ factor formula [101].
@@
 (c) Task 3
-Figure 9: Representation of each task in both isometric (left) and top view (right), along with their respective optimal solutions retrieved by the optimizer.
+Figure 9: Representation of each task in both isometric (left) and top view (right), along with their respective solutions.
 4.2 Tasks
@@
 9(c)).
-We executed each task with a maximum number of nodes $n=20$, bounds on joint rotation set to $\Delta_{\theta}=\pi/4$ on each direction (a limited value emulating hardware constraints), and bounds on link lengths set to $\Delta_{l}\in[25,70]$ to fit the workspace.
-We repeated each task fifty times for each condition to test all the combinations of parameters discussed in Section 4.1 and summarized in the following Section 4.3.
+We executed each task with a maximum link number $n=20$, joint rotation bounds set to $\Delta_{\theta}=\pi/4$ on each direction (a limited value emulating the physical robot’s capabilities), and link length bounds set to $\Delta_{l}\in[25,70]$ to fit the workspace.
+We repeated each task fifty times.
+Concerning dimensionality, each rotational joint comprises two decision variables (i.e., rotation on $x$ and $y$) multiplied for each $|t|$ target (i.e., each configuration of the robot aims at one specific target in the environment).
+Configurations share link lengths, so they add a fixed constant value $n$.
+The overall dimensionality is $2\cdot|t|\cdot n+n$, which is upper-bounded by $2\cdot|t|\cdot n^{\Delta_{\theta}}+n^{\Delta_{l}}$.
+Specifically, the tasks we tested have the following dimensionalities:
+  • 1.
+    Task 1: $2\cdot 20\cdot 2+20=100$ decision variables;
+  • 2.
+    Task 2: $2\cdot 20\cdot 4+20=180$ decision variables; and
+  • 3.
+    Task 3: $2\cdot 20\cdot 8+20=340$ decision variables.
 4.3 Parameter Combinations
@@
     • (b)
-      27 combinations for PSO: 3 values for inertia weight, 3 cognitive constants, and 3 social constants (see Section 4.1.2 and Table 3);
+      27 combinations for PSO: 3 values for inertia weight, 3 cognitive constants, and 3 social constants (see Section 4.1.2 and Table LABEL:tab:param_pso);
     • (c)
@@
 Since we are dealing with a five-objective optimization problem (three objectives in which one of them is further divided into two, as described in Section 3.2), we used the rank from Rank Partitioning as the evaluation metric.
-The Rank Partitioning algorithm, described in Section 3.4, sorts the retrieved solutions based on their objective values according to their predefined priorities; then, it assigns a rank to each of them that linearly increments in ascending order from best to worst – i.e., the best solution gets rank $=1$, the second best gets rank $=2$, and so on.
+The Rank Partitioning algorithm, described in Section 2 of the Supplementary Material, sorts the retrieved solutions based on their objective values according to their predefined priorities; then, it assigns a rank to each of them that linearly increments in ascending order from best to worst – i.e., the best solution gets rank $=1$, the second best gets rank $=2$, and so on.
 Therefore, we evaluated the results by collecting the $2900$ retrieved solutions for each task and assigned them a rank with Rank Partitioning – i.e., from $1$ to $2900$.
@@
 This is interesting, considering that BBBC is a less-known metaheuristic, which yet could compete with big names in the state-of-the-art.
-In conclusion, we found that our optimizer can retrieve high-performance 3D designs in terms of precision in reaching targets, resource consumption, and actuation, especially when the core EC algorithm is GA.
-4.7 Considerations on Problem Dimensionality
-The tasks described in Section 4.2 are optimization problems with high dimensionality in the variable search space.
-This can be quantified from the fixed value of the maximum link number $n=20$ and the number of targets $|t|$ in the task.
-Each link accounts for two decision variables per joint (i.e., rotation on $x$ and $y$), which is multiplied for each desired target (i.e., each configuration of the robot aims at one specific target in the environment.
-The link lengths are shared among all configurations, so they add a constant term.
-This results in a dimensionality of $2\cdot n\cdot|t|+n$.
-We can then evaluate the following dimensionalities:
-  • 1.
-    Task 1 (Fig. 9(a)): $2\cdot 20\cdot 2+20=100$ decision variables;
-  • 2.
-    Task 2 (Fig. 9(b)): $2\cdot 20\cdot 4+20=180$ decision variables; and
-  • 3.
-    Task 3 (Fig. 9(c)): $2\cdot 20\cdot 8+20=340$ decision variables.
+Lastly, we discuss the potential of the optimizer to solve problems with high dimensionality – with reference to the values reported in Section 4.2.
 In the previous 2D work [67], the algorithm was tested only up to 140 decision variables.
 The tasks analyzed in the current work are inherently more complex due to the third dimension of the environment; yet, as shown by the results reported in Section 4.5, the optimizer can tackle high dimensionality problems while still achieving (sub)optimal designs – especially when the core EC algorithm is GA.
-It is reasonable to claim that the limit on dimensionality depends on the computational cost of the specific task.
+In conclusion, we found that our optimizer can retrieve high-performance 3D designs in terms of precision in reaching targets, resource consumption, and actuation, especially when the core EC algorithm is GA.
 5 Conclusion
-In this work, we presented an extension of our innovative tool to optimize the design of soft growing robots for solving specific tasks.
+This work presents an extension of our tool for soft-growing design optimization [67], offered to robotic enthusiasts when facing a specific task.
 While the previous work was limited only to planar manipulators (i.e., in 2D), the current work completes the tool addressing more realistic 3D scenarios.
@@
 Supplementary Information
-A document containing the supplementary material can be downloaded at https://www.mediafire.com/file/s4r3yqm489cccg6/3D+SGR+supplementary+material.pdf/file.
+A document containing the Supplementary Material can be downloaded at https://www.mediafire.com/file/s4r3yqm489cccg6/3D_SGR_supplementary_material.pdf/file.
 A graphical presentation on the Rank Partitioning algorithm, including theory and hands calculation, is available at https://www.mediafire.com/file/6jlp1yr4p8yjjul/22._Rank_Partitioning_for_MOEAs.pdf/file.